### v1

# Stable degeneration of families of klt singularities with constant local volume

Abstract.
We prove that for a locally stable family of klt singularities with constant local volume, the ideal sequences of the minimizing valuations for the normalized volume function form a family of ideals with flat cosupport, which induces a degeneration to a locally stable family of K-semistable log Fano cone singularities. Our proof is a family version of the method of C. Xu and Z. Zhuang proving finite generation by Kollár models and multiple degenerations.
1. Introduction
The study of normalized volumes of real valuations on a klt singularity $x\in(X,\Delta)$ and its minimizer was initiated in [Li_vol], focused on the Stable Degeneration Conjecture proposed in [Li_vol, Conj. 7.1] and [LX_stability, Conj. 1.2]. Following standard notations, $\widehat{\mathrm{vol}}_{X,\Delta}(v)\coloneqq A_{X,\Delta}(v)^{\dim X}\cdot\mathrm{vol}_{X}(v)$ is the normalized volume of a real valuation $v\in\mathrm{Val}_{X,x}$, and $\widehat{\mathrm{vol}}(x;X,\Delta)\coloneqq\inf_{v}\widehat{\mathrm{vol}}_{X,\Delta}(v)$ is the local volume of the klt singularity.
This Stable Degeneration Conjecture has been proved in full generality: There exists a valuation $v$ minimizing $\widehat{\mathrm{vol}}_{X,\Delta}$, which is unique up to scaling and is quasi-monomial; the associated graded algebra $R_{0}=\mathrm{gr}_{v}(\mathscr{O}_{X,x})$ is finitely generated, and induces a klt degeneration $(X_{0}=\mathop{\mathrm{Spec}}(R_{0}),\Delta_{0})$, which is a K-semistable log Fano cone singularity under the natural torus action, and is the unique such K-semistable degeneration. See [LLX_guide] and [Zhuang_survey] for surveys and more comprehensive references.
The uniqueness suggests that the minimizing valuation $v$, or rather the K-semistable degeneration, is associated with a klt singularity canonically. Then it is natural to expect that they should form a family, in an appropriate sense, for a family of klt singularities. As the local volume function $\widehat{\mathrm{vol}}$ is constructible and lower semi-continuous by [Xu_qm] and [BL_lsc], we consider families with constant local volume.
Instead of the minimizing valuation $v$ itself, we will consider the ideal sequence $\mathfrak{a}_{\bullet}=\{\mathfrak{a}_{\lambda}\}_{\lambda}$ associated with it, where $\mathfrak{a}_{\lambda}\coloneqq\{f:v(f)\geq\lambda\}$. Note that the ideal sequence occurs in the definition of the volume and the associated graded algebra. We will prove the following:
Theorem 1.1 (Corollary 5.2).
Let $S$ be a semi-normal scheme essentially of finite type over a field of characteristic zero. Let $\pi\colon x\in(X,\Delta)\to S$ be a locally stable family of klt singularities such that
$s\mapsto\widehat{\mathrm{vol}}(x_{s};X_{s},\Delta_{s})$
is a locally constant function on $S$. Suppose $v^{\mathrm{m}}_{s}\in\mathrm{Val}_{X_{s},x_{s}}$ is a minimizer of the normalized volume for $x_{s}\in(X_{s},\Delta_{s})$, scaled such that $A_{X_{s},\Delta_{s}}(v^{\mathrm{m}}_{s})=1$ for all $s\in S$. Then there is an ideal sequence $\mathfrak{a}_{\bullet}\subseteq\mathscr{O}_{X}$ centered at $x$ such that the following hold:
  1. (1)
    $\mathfrak{a}_{\lambda}/\mathfrak{a}_{>\lambda}$ is flat over $S$ for all $\lambda\geq 0$.
  2. (2)
    Let
    $X_{0}=\mathop{\mathrm{Spec}_{S}}\bigoplus_{\lambda\geq 0}\mathfrak{a}_{\lambda}/\mathfrak{a}_{>\lambda}$
    with the canonical section $x_{0}\in X_{0}(S)$ and the induced divisor $\Delta_{0}$, then $x_{0}\in(X_{0},\Delta_{0})\to S$ is a locally stable family of K-semistable log Fano cone singularities.
  3. (3)
    For every $s\in S$, $\mathfrak{a}_{s,\bullet}$ is the ideal sequence associated with $v_{s}^{\mathrm{m}}$, and the fiber $x_{0,s}\in(X_{0,s},\Delta_{0,s})$ is the degeneration to a K-semistable log Fano cone singularity induced by $v_{s}^{\mathrm{m}}$.
We note that a similar result is recently announced in [Odaka_moduli, Thm. 3.6].
Our proof uses the theory of Kollár models developed in [XZ_stable_deg], which is a higher rank analogue of Kollár components. In [XZ_stable_deg], it was shown that the minimizing valuation $v$ occurs on a Kollár model, and every such quasi-monomial valuation has finitely generated associated graded algebra which induces a klt degeneration. We will generalize both steps to families of klt singularities. The key is to construct a family of Kollár models which accommodates the minimizing valuation on each fiber, with the same coordinates when the dual complexes of different fibers are identified:
Theorem 1.2 (Theorem 5.1).
Let $S=\mathop{\mathrm{Spec}}(A)$, where $A$ is a DVR essentially of finite type over a field of characteristic zero, with the generic point $\eta\in S$ and the closed point $s\in S$. Let $\pi\colon x\in(X,\Delta)\to S$ be a locally stable family of klt singularities with
$\widehat{\mathrm{vol}}(x_{\eta};X_{\eta},\Delta_{\eta})=\widehat{\mathrm{vol}}(x_{s};X_{s},\Delta_{s}).$
Suppose $v^{\mathrm{m}}_{\eta}\in\mathrm{Val}_{X_{\eta},x_{\eta}}$ and $v^{\mathrm{m}}_{s}\in\mathrm{Val}_{X_{s},x_{s}}$ are minimizers of the normalized volume for $x_{\eta}\in(X_{\eta},\Delta_{\eta})$ and $x_{s}\in(X_{s},\Delta_{s})$, respectively, scaled such that $A_{X_{\eta},\Delta_{\eta}}(v^{\mathrm{m}}_{\eta})=A_{X_{s},\Delta_{s}}(v^{\mathrm{m}}_{s})$. Then there exists a locally stable family of Kollár models $f\colon(Y,E)\to(X,\Delta)$ at $x$ over $S$ such that
$v^{\mathrm{m}}_{\eta}\in\mathrm{QM}(Y_{\eta},E_{\eta})\quad\text{and}\quad v^{\mathrm{m}}_{s}\in\mathrm{QM}(Y_{s},E_{s}),$
and they are identified under the canonical isomorphism $\mathrm{QM}(Y_{\eta},E_{\eta})\simeq\mathrm{QM}(Y_{s},E_{s})$.
Theorem 1.1 implies a representability for semi-normal schemes:
Theorem 1.3 (Corollary 5.4).
[…]
Outline of the proof
Suppose we are in the situation of the Theorems above, where the base $S$ is the spectrum of a DVR. It is easy to degenerate the ideal sequence $\mathfrak{a}_{\bullet}(v^{\mathrm{m}}_{\eta})$ on $X_{\eta}$ to an ideal sequence $I_{\bullet}$ on $X$ with flat cosupport over $S$. Then $I_{\bullet}$ has constant multiplicity over $S$; by the lower semi-continuity of the lc thresholds, the closed fiber $I_{s,\bullet}$ computes $\widehat{\mathrm{vol}}(x_{s};X_{s},\Delta_{s})$. By a stronger uniqueness of minimizer implicitly in [BLQ_filt], we conclude that $\mathfrak{a}_{s,\bullet}$ is the saturation of $I_{s,\bullet}$; see Lemma 2.17. We collect some other preliminaries, and also fix relevant notations, in Section 2.
This leads us to consider a more general situation where $I_{\bullet}$ is a family of ideal sequences on $X$ with constant lc thresholds, and $v_{\eta}$ computes the lc threshold of $I_{\eta,\bullet}$ on $X_{\eta}$. Then we can construct a model $f\colon(Y,E)\to(X,\Delta)$ of Fano type such that $v_{\eta}\in\mathrm{QM}(Y_{\eta},E_{\eta})$, by a similar argument that has appeared in [LX_stability, Thm. 4.2], [Xu_K-Stability_Book, Thm. 4.48], and [XZ_stable_deg, Lem. 3.2]. In our case, the assumption that $I_{\bullet}$ has constant lc threshold implies that $\pi\circ f\colon(Y,f^{-1}_{*}\Delta+E)\to S$ is a locally stable family, hence we can find $v_{s}\in\mathrm{QM}_{z}(Y_{s},E_{s})$, where $z$ is a generic point of $\overline{\{\zeta\}}\cap Y_{s}$, that computes the lc threshold of $I_{s,\bullet}$. If $I_{\bullet}$ is the ideal sequence computing $\widehat{\mathrm{vol}}$ as in the last paragraph, then $v_{s}=v^{\mathrm{m}}_{s}$.
Now the closed fiber $(Y_{s},E_{s})$ is lc by the construction, whereas we would like it to be qdlt, so that we can get a family of Kollár models. The key is that $v^{\mathrm{m}}_{\eta}$ and $v^{\mathrm{m}}_{s}$ are special valuations, namely, monomial lc places of special $\mathbb{Q}$-complements, so they compute the lc threshold of $I_{\bullet}$ after any small perturbation of $\Delta$ to $\Delta+\epsilon D$. The heuristic is that such perturbations exclude any lc center on $(Y_{s},E_{s})$ that does not contain $z$. This method was used in the proof of [XZ_stable_deg, Prop. 4.6]. In out situation, however, the model $(Y,E)$ constructed above may change when we perturb the boundary.
Instead, note that $v^{\mathrm{m}}_{s}$ can be viewed as the limit of the weighted blow-up valuations of $v^{\mathrm{m}}_{\eta}$ and the discrete valuation $v_{0}$ induced by $S$, when the weight of $v_{0}$ goes to +$\infty$+ ∞. Then the perturbation allows us to find a special $\mathbb{Q}$-complement for which all these weighted blow-ups are monomial lc places. Hence we get a model of qdlt Fano type model accommodating them by [XZ_stable_deg, Thm. 3.14], whose anti-canonical model is a locally stable family of Kollár models as desired.
In Section 3, we study models of qdlt Fano type and special valuations; we extend some results in [XZ_stable_deg], and make the perturbation property above precise in Lemma 3.8. The constructions of families of models are in Section 4, where all the results are stated over a regular local ring, since the proofs are exactly the same as in the DVR case. We also prove the multiple equivariant degeneration in [Xu_towards] and [XZ_stable_deg, §4.2] for locally stable families. Finally, we prove the main theorems in Section 5.
Acknowledgement
I would like to thank my advisor Chenyang Xu for suggesting me this topic, and for constant support and guidance on this research. I would also like to thank my cohorts Junyao Peng and Lu Qi for inspiring discussions. The research is partly supported by a Simons Foundation grant.
2. Preliminaries
2.1. Birational geometry and singularities
All schemes that we consider will be quasi-compact, separated and essentially of finite type over a field of characteristic zero.
A pair $(X,\Delta)$ consists of a reduced, equidimensional scheme $X$ with a Weil $\mathbb{Q}$-divisor $\Delta$ on $X$. It is said to be local if $X$ is the spectrum of a local ring. A singularity $x\in(X,\Delta)$ consists of an affine pair $(X,\Delta)$ essentially of finite type over a field $k$ of characteristic 00, and a $k$-point $x\in X$ [^1] It is possible to consider arbitrary closed points, and on more general excellent schemes. But many theorems for the normalized volume are only proved in this case in the literature, sometimes even assuming $k$ is algebraically closed. Also, requiring $x\in X(k)$ is compatible with our definition that a family of singularities is given by a section.
We follow the standard terminology about singularities of pairs in [Kollar_singularity] when $\Delta\geq 0$, but add the prefix “sub-” when $\Delta$ is not necessarily effective, e.g., sub-lc, sub-klt.
2.1.1. Qdlt pairs
We recall the concept of qdlt pairs, following [dFKX].
Definition 2.1 ([dFKX, Def. 35], [XZ_stable_deg, Def. 2.1, 2.2]).
A pair $(X,\Delta)$ is said to be simple-toroidal if $\Delta$ is reduced and every $x\in X$ has an open neighborhood $U\subseteq X$ such that $(U,\Delta|_{U})\simeq(V,D)/G$, where $(V,D)$ is an snc pair, and $G$ is a finite abelian group acting on $V$ that preserves each component of $D$ and acts freely on $V\smallsetminus D$.
A pair $(X,\Delta)$ is said to be qdlt if it is lc, and $(\mathop{\mathrm{Spec}}\mathscr{O}_{X,z},\Delta|_{\mathop{\mathrm{Spec}}\mathscr{O}_{X,z}})$ is simple-toroidal for every lc center $z\in(X,\Delta)$.
Lemma 2.2 ([dFKX, Prop. 34]).
Let $(X,\Delta)$ be a local slc pair of dimension $r$. Suppose $D_{1},\ldots,D_{r}$ are reduced $\mathbb{Q}$-Cartier divisors on $X$ such that $D_{1}+\cdots+D_{r}\leq\Delta$. Then $(X,\Delta)$ is simple-toroidal, and $\Delta=D_{1}+\cdots+D_{r}$, where $D_{i}$ is normal and irreducible.
2.1.2. Models
Definition 2.3.
Let $(X,\Delta)$ be a pair. A model
$f\colon(Y,E)\to(X,\Delta)$
for $(X,\Delta)$ consists of a pair $(Y,E)$ where $Y$ is a normal scheme and $E$ is a reduced Weil divisor on $Y$, and a projective birational morphism $f\colon Y\to X$. It is called a model at a point $x\in X$ if all irreducible components of $E$ are centered at $x$.
A model $f\colon(Y,E)\to(X,\Delta)$ is said to be snc (resp., simple-toroidal) if $\mathrm{Ex}(f)$ has pure codimension one in $Y$, and $(Y,\mathrm{Supp}(f^{-1}_{*}\Delta+\mathrm{Ex}(f)+E))$ is snc (resp., simple-toroidal). It is called a log resolution [^2] This seems to be a non-standard use of terminology as we allow log resolutions to be only toroidal, rather than snc. However, the theory of resolution of singularities works as usual. For example, we can take a log resolution that is a local isomorphism over the simple-toroidal locus (see [ATW_log_resolution]). Anyway, toroidal pairs are logarithmic regular in the sense of logarithmic geometry. if $E$ contains all components of $f^{-1}_{*}\Delta$ and $\mathrm{Ex}(f)$.
Lemma 2.4.
Let $(X,\Delta)$ be a klt pair, and $\mathfrak{a}\subseteq\mathscr{O}_{X}$ be a coherent ideal with $\mathrm{lct}(X,\Delta;\mathfrak{a}^{c})\geq 1$ for some $c>0$. Suppose $E_{1},\ldots,E_{r}$ are exceptional prime divisors over $X$ such that
$A_{X,\Delta+\mathfrak{a}^{c}}(E_{i})\coloneqq A_{X,\Delta}(E_{i})-c\mathop{\mathrm{ord}_{E_{i}}}(\mathfrak{a})<1$
for all $i=1,\ldots,r$. Then there exists a model $f\colon(Y,E)\to(X,\Delta)$ where $E=\sum_{i=1}^{r}E_{i}$ is the sum of all $f$-exceptional divisors. Moreover, we may assume either of the following holds:
  1. (i)
    [BCHM, Cor. 1.4.3] $Y$ is $\mathbb{Q}$-factorial; or
  2. (ii)
    [Mor_extract, Thm. 1] $\mathrm{Ex}(f)=E$ has pure codimension $1$, each $E_{i}$ is $\mathbb{Q}$-Cartier, and there exists an $f$-ample divisor $-A$ on $Y$ with $A\geq 0$ and $\mathrm{Supp}(A)=E$.
2.1.3. Locally stable families
We use the notion of locally stable families of pairs following [Kol_fam].
Let $S$ be a reduced scheme. A family of pairs $\pi\colon(X,\Delta)\to S$ is defined in [Kol_fam, Def. 4.3], and local stability of $\pi\colon(X,\Delta)\to S$ is defined in [Kol_fam, Def. 4.7].
Definition 2.5.
Let $S$ be a reduced scheme. A family of singularities $\pi\colon x\in(X,\Delta)\to S$ consists of a family of pairs $\pi\colon(X,\Delta)\to S$ where $\pi$ is affine, together with a section $x\colon S\to X$. Thus in each fiber $X_{s}$ at a point $s\in S$, the section gives a closed point $x_{s}\in X_{s}$. The closed subscheme $x(S)\subseteq X$ is also denoted by $\{x\}$.
Definition 2.6.
Let $S$ be a reduced scheme, $\pi\colon(X,\Delta)\to S$ be a family of pairs. A family of models
$f\colon(Y,E)\to(X,\Delta)$
consists of a projective birational morphism $f\colon Y\to X$ such that $\pi\circ f$ is flat, and a relative Mumford divisor $E$ on $Y/S$ (see [Kol_fam, Def. 4.68]), such that every fiber $f_{s}\colon(Y_{s},E_{s})\to(X_{s},\Delta_{s})$ is a model.
Let $\pi\colon x\in(X,\Delta)\to S$ be a family of singularities. Then $f\colon(Y,E)\to(X,\Delta)$ is called a family of models at $x$ if moreover every fiber $f_{s}\colon(Y_{s},E_{s})\to(X_{s},\Delta_{s})$ is a model at $x_{s}$.
2.2. Valuations
2.2.1. Real valuations
All valuations that we consider will be real valuations.
Let $X$ be an integral scheme with function field $K(X)$. A real valuation $v$ on $K(X)$ gives a valuation ring $\mathcal{O}_{v}$ of $K(X)$ of rank $1$ with an order-preserving embedding of the valuation group $\Gamma_{v}\hookrightarrow\mathbb{R}$, which recovers the valuation $v\colon K(X)^{\times}\to\mathbb{R}$. By convention, we set $v(0)=+\infty$.
The rational rank of $v\in\mathrm{Val}_{X}$ is $\mathop{\mathrm{rat.rank}}(v)\coloneqq\dim_{\mathbb{Q}}(\Gamma_{v}\otimes_{\mathbb{Z}}\mathbb{Q})$.
The set of all real valuations on $K$ centered on $X$ is denoted by $\mathrm{Val}_{X}$. For $x\in X$, let
$\mathrm{Val}_{X,x}\coloneqq\{v\in\mathrm{Val}_{X}:\mathrm{center}_{X}(v)=x\}.$
More generally, if $X=\bigcup_{i}X_{i}$ is a reduced scheme, where $X_{i}$ are the irreducible components of $X$, then we write $\mathrm{Val}_{X}\coloneqq\bigsqcup_{i}\mathrm{Val}_{X_{i}}$.
Suppose $f\colon Y\to X$ is a dominant morphism, where $Y$ is an integral scheme, then we have an induced map $\mathrm{Val}_{Y}\to\mathrm{Val}_{X}$ by restricting a valuation on $K(Y)$ to the subfield $K(X)$ via $f$. If $f$ is birational and proper, then this is a bijection, and we usually identify $\mathrm{Val}_{Y}=\mathrm{Val}_{X}$.
A real valuation $v\in\mathrm{Val}_{X}$ can be evaluated at functions, coherent ideals, and Cartier divisors on $X$, as well as their $\mathbb{R}$-linear combinations (e.g., $\mathbb{R}$-Cartier $\mathbb{R}$-divisors).
2.2.2. Quasi-monomial valuations
Let $(Y,E)$ be an snc pair. Suppose $y$ is a generic point of a stratum $E_{1}\cap\cdots\cap E_{r}\subseteq Y$, and $t_{1},\ldots,t_{r}\in\mathscr{O}_{Y,y}$ is a regular system of parameters such that $E_{i}$ is defined by $t_{i}$ near $y$ for all $i=1,\ldots,r$. Then there is an isomorphism $\hat{\mathscr{O}}_{Y,y}\simeq\kappa(y)[\![t_{1},\ldots,t_{r}]\!]$.
For $\alpha=(\alpha_{1},\ldots,\alpha_{r})\in\mathbb{R}_{\geq 0}^{r}$, there exists a unique real valuation $v_{\alpha}\in\mathrm{Val}_{Y}$ satisfying the following: if $f\in\mathscr{O}_{Y,y}$ is written as $f=\sum_{m\in\mathbb{N}^{r}}c_{m}t_{1}^{m_{1}}\cdots t_{r}^{m_{r}}$ in $\hat{\mathscr{O}}_{Y,y}\simeq\kappa(y)[\![t_{1},\ldots,t_{r}]\!]$, then
$v_{\alpha}(f)=\inf\{\langle\alpha,m\rangle:c_{m}\neq 0\}\in\mathbb{R}_{\geq 0}\cup\{+\infty\},\quad\text{where}\ \langle\alpha,m\rangle=\sum_{i=1}^{r}\alpha_{i}m_{i};$
see [JM, Prop. 3.1]. We call $v_{\alpha}$ a monomial valuation at $y\in(Y,E)$, and denote by $\mathrm{QM}_{y}(Y,E)$ the set of all monomial valuations at $y$. Finally, we define $\mathrm{QM}(Y,E)\coloneqq\bigcup_{y}\mathrm{QM}_{y}(Y,E)$, where $y$ ranges over all generic points of all strata of $(Y,E)$.
Following [XZ_stable_deg], we consider a generalization of monomial valuations to simple-toroidal pairs.
Definition 2.7.
Let $(Y,E=\sum_{i}E_{i})$ be a simple-toroidal pair. Suppose $y\in Y$ is a generic point of a stratum $\bigcap_{i=1}^{r}E_{i}\subseteq Y$. Thus locally at $y$ there is a finite abelian cover
$\mu\colon(Y^{\prime},E^{\prime})\to(Y,E)$
which is étale away from $E$, such that $(Y^{\prime},E^{\prime}=\sum_{i}E^{\prime}_{i})$ is snc, and $\mu^{*}E_{i}=m_{i}E_{i}^{\prime}$ for some $m_{i}>0$. Let $y^{\prime}\in\mu^{-1}(y)$, then $y^{\prime}$ is a generic point of $\bigcap_{i=1}^{r}E_{i}^{\prime}\subseteq Y^{\prime}$. For $\alpha=(\alpha_{1},\ldots,\alpha_{r})\in\mathbb{R}_{\geq 0}^{r}$, define $v_{\alpha}\in\mathrm{Val}_{Y}$ to be the restriction of $v^{\prime}_{\alpha^{\prime}}\in\mathrm{QM}_{y^{\prime}}(Y^{\prime},E^{\prime})$ corresponding to $\alpha^{\prime}=(\alpha_{1}/m_{1},\ldots,\alpha_{r}/m_{r})$. Thus
$v_{\alpha}(E_{i})=v^{\prime}_{\alpha^{\prime}}(\mu^{*}E_{i})=m_{i}v^{\prime}_{\alpha^{\prime}}(E_{i}^{\prime})=\alpha_{i}.$
Note that $v_{\alpha}$ does not depend on the choice of a finite abelian cover.
We call $v_{\alpha}$ a monomial valuation at $y\in(Y,E)$, and denote by $\mathrm{QM}_{y}(Y,E)\subseteq\mathrm{Val}_{Y}$ the set of all such $v_{\alpha}$. The relative interior $\mathrm{QM}_{y}^{\circ}(Y,E)\subseteq\mathrm{QM}_{y}(Y,E)$ consists of $v_{\alpha}$ for which $\alpha\in\mathbb{R}_{>0}^{r}$. By definition there is a canonical bijection
$\mathrm{QM}_{y}(Y,E)\simeq\mathbb{R}_{\geq 0}^{r},\quad v\mapsto(v(E_{1}),\ldots,v(E_{r})),$
which is in fact a homeomorphism when $\mathrm{Val}_{Y}$ is equipped with a suitable topology.
More generally, let $(Y,E)$ be a pair where $E=\sum_{i}E_{i}$ is reduced, such that $(\mathop{\mathrm{Spec}}\mathscr{O}_{Y,y},E|_{\mathrm{Spec}\mathscr{O}_{Y,y}})$ is simple-toroidal for all generic points $y$ of all strata $\bigcap_{i\in I}E_{i}\subseteq Y$. Then we define
$\mathrm{QM}(Y,E)\coloneqq\bigcup\nolimits_{y}\mathrm{QM}_{y}(Y,E)=\bigsqcup\nolimits_{y}\mathrm{QM}^{\circ}_{y}(Y,E)$
where $\mathrm{QM}_{y}(Y,E)\coloneqq\mathrm{QM}_{y}(\mathop{\mathrm{Spec}}\mathscr{O}_{Y,y},E|_{\mathrm{Spec}\mathscr{O}_{Y,y}})$.
Remark.
Let $(Y,E)$ be a pair where $E$ is a reduced Weil divisor. Then $(Y,E)$ is simple-toroidal if and only if $Y\smallsetminus E\hookrightarrow Y$ is a toroidal embedding without self-intersection (in the sense of [KKMSD_toroidal, §II.1]) such that the rational convex polyhedral cones $\sigma^{Z}\subseteq N^{Z}_{\mathbb{R}}$ associated to all strata $Z\subseteq Y$ are simplicial. If $z$ is the generic point of $Z$, then there is a canonical isomorphism
$\sigma^{Z}\simeq\mathrm{QM}_{z}(Y,E),$
and $\sigma^{Z}\cap N^{Z}$ corresponds to the lattice generated by valuations given by prime divisors over $Y$. Hence we can equip $\mathrm{QM}(Y,E)$ with an integral structure.
If $f\colon(Y^{\prime},E^{\prime})\to(Y,E)$ is a toroidal resolution such that $(Y^{\prime},E^{\prime})$ is snc, then $\mathrm{QM}(Y^{\prime},E^{\prime})=\mathrm{QM}(Y,E)$. If $y^{\prime}\in(Y^{\prime},E^{\prime})$ is a generic point of a stratum, then $\mathrm{QM}_{y^{\prime}}(Y^{\prime},E^{\prime})\subseteq\mathrm{QM}_{f(y)}(Y,E)$ is a rational simplicial cone generated by an integral basis.
Conversely, suppose $\sigma\subseteq\mathrm{QM}_{y}(Y,E)$ is a rational simplicial cone, then there exists a toroidal proper birational morphism $f\colon(Y^{\prime},E^{\prime})\to(Y,E)$ such that $(Y^{\prime},E^{\prime})$ is simple-toroidal and $\mathrm{QM}_{y^{\prime}}(Y^{\prime},E^{\prime})=\sigma$ for some $y^{\prime}\in f^{-1}(y)$; furthermore, one may assume that every $f$-exceptional divisor corresponds to an edge of $\sigma$, and $\mathrm{Ex}(f)$ supports an $f$-ample divisor (see [Xu_K-Stability_Book, Lem. 5.28]).
Lemma 2.8.
Let $(Y,E+F)$ be a local simple-toroidal pair, where $E=\sum_{i=1}^{r}E_{i}$ and $F=\sum_{j=1}^{s}F_{j}$ have no common components. Let $Y_{0}=\bigcap_{j=1}^{s}F_{j}$, and $E_{0}=E|_{Y_{0}}$. Then $(Y_{0},E_{0})$ is simple-toroidal, and there are canonical maps
[…]
such that if $v_{\alpha}\in\mathrm{QM}(Y,E)$ maps to $\bar{v}_{\alpha}\in\mathrm{QM}(Y_{0},E_{0})$, and $\beta=(\beta_{j})\in\mathbb{R}_{\geq 0}^{s}$, then
$v_{\alpha}(f)\leq v_{\alpha,\beta}(f)\leq\sup\{v_{\alpha,\beta}(f):\beta\in\mathbb{R}_{\geq 0}^{s}\}=\lim_{\beta\to\infty}v_{\alpha,\beta}(f)=\bar{v}_{\alpha}(\bar{f}),$
for every $f\in\mathscr{O}_{Y}$, where $\bar{f}\in\mathscr{O}_{Y_{0}}$ is the image of $f$, and $\beta\to\infty$ means each $\beta_{j}\to\infty$.
Proof.
By localization and passing to a finite abelian cover, we may assume that $(Y,E+F)$ is snc, and $\bigcap_{i=1}^{r}E_{i}\cap\bigcap_{j=1}^{s}F_{j}=\{y\}$, where $y\in Y$ is the closed point. Write $R=\mathscr{O}_{Y,y}$, and take a regular system of parameters $t_{1},\ldots,t_{r},x_{1},\ldots,x_{s}\in R$ such that $E_{i}=\mathrm{div}(t_{i})$ and $F_{j}=\mathrm{div}(x_{j})$. Then
$\mathscr{O}_{Y_{0},y}\simeq R_{0}\coloneqq R/(x_{1},\ldots,x_{s})$
is a regular local ring, and $\bar{t}_{1},\ldots,\bar{t}_{r}\in R_{0}$ is a regular system of parameters defining $E_{0}$. Write
$f=\sum_{m\in\mathbb{N}^{r}}\sum_{n\in\mathbb{N}^{s}}c_{m,n}t_{1}^{m_{1}}\cdots t_{r}^{m_{r}}x_{1}^{n_{1}}\cdots x_{s}^{n_{s}}\in\hat{R}\simeq\kappa(y)[\![t_{1},\ldots,t_{r},x_{1},\ldots,x_{s}]\!],$
so that
$\bar{f}=\sum_{m\in\mathbb{N}^{r}}c_{m,0}\bar{t}_{1}^{m_{1}}\cdots\bar{t}_{r}^{m_{r}}\in\hat{R}_{0}\simeq\kappa(y)[\![\bar{t}_{1},\ldots,\bar{t}_{r}]\!].$
Hence
$v_{\alpha}(f)=\inf\{\langle\alpha,m\rangle:c_{m,n}\neq 0\}$ $\leq\inf\{\langle\alpha,m\rangle+\langle\beta,n\rangle:c_{m,n}\neq 0\}$
$\leq\sup_{\beta\geq 0}\inf\{\langle\alpha,m\rangle+\langle\beta,n\rangle:c_{m,n}\neq 0\}$
$=\lim_{\beta\to\infty}\inf\{\langle\alpha,m\rangle+\langle\beta,n\rangle:c_{m,n}\neq 0\}$
$=\inf\{\langle\alpha,m\rangle:c_{m,0}\neq 0\}=\bar{v}_{\alpha}(\bar{f}).$
since $\langle\alpha,m\rangle+\langle\beta,n\rangle=\sum_{i=1}^{r}\alpha_{i}m_{i}+\sum_{j=1}^{s}\beta_{j}n_{j}\geq\min_{j}\beta_{j}$ when $n\neq 0$. ∎
Definition 2.9.
Let $X$ be an integral scheme. A real valuation $v\in\mathrm{Val}_{X}$ is said to be quasi-monomial if $v\in\mathrm{QM}(Y,E)$ for some snc pair $(Y,E)$ with a birational morphism $f\colon Y\to X$ of finite type. The set of quasi-monomial valuations on $X$ is denoted by $\mathrm{Val}_{X}^{\mathrm{qm}}\subseteq\mathrm{Val}_{X}$. It is equivalent to require $(Y,E)$ to be simple-toroidal.
Note that divisorial valuations are exactly quasi-monomial valuations of rational rank $1$. The set of divisorial valuations on $X$ is denoted by $\mathrm{Val}_{X}^{\mathrm{div}}\subseteq\mathrm{Val}_{X}^{\mathrm{qm}}$.
2.2.3. Log discrepancies
Let $(X,\Delta)$ be a pair such that $K_{X}+\Delta$ is $\mathbb{Q}$-Cartier. Suppose $f\colon Y\to X$ is a birational morphism of finite type, and $F$ is a prime divisor on $Y$. Then the log discrepancy
$A_{X,\Delta}(F)\coloneqq 1+\mathrm{ord}_{F}(K_{Y}-f^{*}(K_{X}+\Delta))$
only depends on the valuation $\mathrm{ord}_{F}\in\mathrm{Val}_{X}$. The log discrepancy extends to a function
$A_{X,\Delta}\colon\mathrm{Val}_{X}\to\mathbb{R}\cup\{+\infty\}$
such that
$A_{X,\Delta}(v)\coloneqq\sup_{(Y,E)}\sum\nolimits_{i}v(E_{i})\cdot A_{X,\Delta}(E_{i})$
where the supremum ranges over all log resolutions $f\colon(Y,E=\sum_{i}E_{i})\to(X,\Delta)$; see [JM, §5].
Lemma 2.10.
Let $(X,\Delta)$ be a pair such that $K_{X}+\Delta$ is $\mathbb{Q}$-Cartier. Suppose
$f\colon(Y,E=E_{1}+\cdots+E_{r})\to(X,\Delta)$
is a log resolution. Then
$A_{X,\Delta}(v)\geq\sum_{i=1}^{r}v(E_{i})\cdot A_{X,\Delta}(E_{i})$
for all $v\in\mathrm{Val}_{X}$, and the equality holds if and only if $v\in\mathrm{QM}(Y,E)$.
Proof.
Note that the value $\sum_{i=1}^{r}v(E_{i})\cdot A_{X,\Delta}(E_{i})$ does not change if we replace $(Y,E)$ with a toroidal resolution, and the set $\mathrm{QM}(Y,E)$ also remains the same. Hence we may assume that $f\colon(Y,E)\to(X,\Delta)$ is an snc model, so the inequality follows by definition.
If the equality holds, then it is proved that $v\in\mathrm{QM}(Y,E)$ in [JM, Cor. 5.4] when $X$ is regular and $\Delta=0$. In general, we can reduce it to the regular case as follows. Let $\Delta_{Y}$ be the crepant pullback of $\Delta$, so that
$\Delta_{Y}=\sum_{i=1}^{r}(1-A_{X,\Delta}(E_{i}))\cdot E_{i},$
and
$A_{X,\Delta}(v)=A_{Y,\Delta_{Y}}(v)=A_{Y}(v)-\sum_{i=1}^{r}v(E_{i})\cdot(1-A_{X,\Delta}(E_{i})).$
If $A_{X,\Delta}(v)=\sum_{i=1}^{r}v(E_{i})\cdot A_{X,\Delta}(E_{i})$, then $A_{Y}(v)=\sum_{i=1}^{r}v(E_{i})$, hence $v\in\mathrm{QM}(Y,E)$. ∎
Definition 2.11.
Let $(X,\Delta)$ be an lc pair. A real valuation $v\in\mathrm{Val}_{X}$ is called an lc place for $(X,\Delta)$ if $A_{X,\Delta}(v)=0$. Let $\mathrm{LCP}(X,\Delta)\subseteq\mathrm{Val}_{X}$ denote the set of all lc places for $(X,\Delta)$. A point $z\in X$ is called an lc center of $(X,\Delta)$ if $z=\mathrm{center}_{X}(v)$ for some $v\in\mathrm{LCP}(X,\Delta)$. Note that $\mathrm{LCP}(X,\Delta)\subseteq\mathrm{Val}_{X}^{\mathrm{qm}}$.
2.3. Local volumes
2.3.1. Ideal sequences
We consider ideal sequences indexed by $\mathbb{R}_{\geq 0}$. They are called graded systems of ideals by [ELS], and filtrations by [BLQ_filt].
Definition 2.12.
Let $X$ be a scheme. An ideal sequence $\mathfrak{a}_{\bullet}=\{\mathfrak{a}_{\lambda}\}_{\lambda\geq 0}$ on $X$ consists of coherent ideals $\mathfrak{a}_{\lambda}\subseteq\mathscr{O}_{X}$ for $\lambda\in\mathbb{R}_{\geq 0}$, satisfying the following conditions:
  1. (1)
    $\mathfrak{a}_{\lambda}\subseteq\mathfrak{a}_{\mu}$ if $\lambda\geq\mu$;
  2. (2)
    $\mathfrak{a}_{\lambda}\mathfrak{a}_{\mu}\subseteq\mathfrak{a}_{\lambda+\mu}$ for all $\lambda,\mu\geq 0$.
  3. (3)
    $\mathfrak{a}_{0}=\mathscr{O}_{X}$, and $\mathfrak{a}_{\lambda}=\bigcap_{\lambda^{\prime}<\lambda}\mathfrak{a}_{\lambda^{\prime}}$ for all $\lambda>0$.
Let $x\in X$ be a point, with the prime ideal $\mathfrak{p}_{x}\subseteq\mathscr{O}_{X}$. We say an ideal sequence $\mathfrak{a}_{\bullet}$ is centered at $x$ if $\mathfrak{a}_{\lambda}$ is $\mathfrak{p}_{x}$-primary for all $\lambda>0$. In this case, $\mathfrak{a}_{\bullet}$ is uniquely determined by the corresponding ideal sequence $\mathfrak{a}_{\bullet}\mathscr{O}_{X,x}=\{\mathfrak{a}_{\lambda}\mathscr{O}_{X,x}\}_{\lambda}$ of $\mathscr{O}_{X,x}$; hence we may assume $x$ is a closed point by localization if necessary.
Suppose $X$ in integral and $v\in\mathrm{Val}_{X}$. The ideal sequence $\mathfrak{a}_{\bullet}(v)$ associated with $v$ is defined by
$\mathfrak{a}_{\lambda}(v)\coloneqq\{f\in\mathscr{O}_{X}:v(f)\geq\lambda\}.$
If $\mathrm{center}_{X}(v)=x$, then $\mathfrak{a}_{\bullet}(v)$ is an ideal sequence centered at $x$.
A real valuation $v$ on $X$ can be evaluated at an ideal sequence $\mathfrak{a}_{\bullet}$ by
$v(\mathfrak{a}_{\bullet})=\inf_{\lambda>0}\frac{v(\mathfrak{a}_{\lambda})}{\lambda}=\lim_{\lambda\to\infty}\frac{v(\mathfrak{a}_{\lambda})}{\lambda}\in\mathbb{R}_{\geq 0}\cup\{+\infty\}.$
Note that $v(\mathfrak{a}_{\bullet}(v))=1$.
Definition 2.13.
Let $X$ be an integral scheme, $x\in X$ be a closed point, and $\mathfrak{a}_{\bullet}=\{\mathfrak{a}_{\lambda}\}_{\lambda\geq 0}$ be an ideal sequence centered at $x$. Let $n=\dim_{x}(X)$. The multiplicity of $\mathfrak{a}_{\bullet}$ is
$\mathrm{e}(\mathfrak{a}_{\bullet})=\mathrm{e}_{X}(\mathfrak{a}_{\bullet})\coloneqq\limsup_{\lambda\to\infty}\frac{\mathrm{length}_{\mathscr{O}_{X}}(\mathscr{O}_{X}/\mathfrak{a}_{\lambda})}{\lambda^{n}/n!}.$
The limit superior is actually a limit by [Cut_limit, Thm. 1.1] under our assumptions for schemes.
Suppose $v\in\mathrm{Val}_{X,x}$. The volume of $v$ is
$\mathrm{vol}(v)=\mathrm{vol}_{X}(v)\coloneqq\mathrm{e}_{X}(\mathfrak{a}_{\bullet}(v)),$
where $\mathfrak{a}_{\bullet}(v)\subseteq\mathscr{O}_{X}$ is the ideal sequence associated with $v$. The set of valuations with positive volume is denoted by $\mathrm{Val}_{X,x}^{+}$. Note that $\mathrm{Val}_{X,x}^{\mathrm{qm}}\subseteq\mathrm{Val}_{X,x}^{+}$ by Izumi’s Theorem; see [BLQ_filt, Prop. 2.7].
Definition 2.14 ([BLQ_filt, Def. 3.1]).
Let $X$ be an integral scheme, $x\in X$ be a closed point, and $\mathfrak{a}_{\bullet}$ be an ideal sequence centered at $x$. The saturation $\tilde{\mathfrak{a}}_{\bullet}$ of $\mathfrak{a}_{\bullet}$ is the ideal sequence defined by
$\tilde{\mathfrak{a}}_{\lambda}=\mathopen{}\mathclose{{}\left\{f\in\mathscr{O}_{X}:v(f)\geq\lambda\cdot v(\mathfrak{a}_{\bullet})\ \text{for all}\ v\in\mathrm{Val}_{X,x}^{\mathrm{div}}}\right\}.$
It has following basic properties:
  1. (1)
    [BLQ_filt, Lem. 3.8] $\mathfrak{a}_{\lambda}\subseteq\tilde{\mathfrak{a}}_{\lambda}$ for all $\lambda\geq 0$;
  2. (2)
    [BLQ_filt, Cor. 3.16] $\mathrm{e}_{X}(\mathfrak{a}_{\bullet})=\mathrm{e}_{X}(\tilde{\mathfrak{a}}_{\bullet})$;
  3. (3)
    [BLQ_filt, Prop. 3.12] $v(\mathfrak{a}_{\bullet})=v(\tilde{\mathfrak{a}}_{\bullet})$ for all $v\in\mathrm{Val}_{X,x}^{+}$.
Lemma 2.15.
Let $x\in(X,\Delta)$ be a klt singularity, and $\mathfrak{a}_{\bullet}$ be an ideal sequence centered at $x$ with
$\mathrm{lct}(X,\Delta;\mathfrak{a}_{\bullet})<+\infty.$
Suppose $\tilde{\mathfrak{a}}_{\bullet}$ is the saturation of $\mathfrak{a}_{\bullet}$. Then
$\mathrm{lct}(X,\Delta;\tilde{\mathfrak{a}}_{\bullet})=\mathrm{lct}(X,\Delta;\mathfrak{a}_{\bullet}),$
and a valuation $v\in\mathrm{Val}_{X,x}$ computes the lc threshold of $\tilde{\mathfrak{a}}_{\bullet}$ if and only if it computes that of $\mathfrak{a}_{\bullet}$.
Proof.
We have
$\mathrm{lct}(X,\Delta;\tilde{\mathfrak{a}}_{\bullet})=\inf_{w\in\mathrm{Val}_{X,x}^{\mathrm{div}}}\frac{A_{X,\Delta}(w)}{w(\tilde{\mathfrak{a}}_{\bullet})}=\inf_{w\in\mathrm{Val}_{X,x}^{\mathrm{div}}}\frac{A_{X,\Delta}(w)}{w(\mathfrak{a}_{\bullet})}=\mathrm{lct}(X,\Delta;\mathfrak{a}_{\bullet})$
since $w(\tilde{\mathfrak{a}}_{\bullet})=w(\mathfrak{a}_{\bullet})$ for every $w\in\mathrm{Val}_{X,x}^{\mathrm{div}}$. For a general real valuation $v\in\mathrm{Val}_{X,x}$, we have
$\frac{A_{X,\Delta}(v)}{v(\tilde{\mathfrak{a}}_{\bullet})}\geq\frac{A_{X,\Delta}(v)}{v(\mathfrak{a}_{\bullet})}$
since $\mathfrak{a}_{\bullet}\subseteq\tilde{\mathfrak{a}}_{\bullet}$. Thus, if $v$ computes the lc threshold of $\tilde{\mathfrak{a}}_{\bullet}$, then it also computes that of $\mathfrak{a}_{\bullet}$. Conversely, assume $v$ computes the lc threshold of $\mathfrak{a}_{\bullet}$. Then $v$ also computes the lc threshold of the ideal sequence $\mathfrak{a}_{\bullet}(v)$ associated with itself by [JM, Thm. 7.8]. Hence
$\inf_{w\in\mathrm{Val}_{X,x}^{\mathrm{div}}}\frac{A_{X,\Delta}(w)}{w(\mathfrak{a}_{\bullet}(v))}=\mathrm{lct}(X,\Delta;\mathfrak{a}_{\bullet}(v))=\frac{A_{X,\Delta}(v)}{v(\mathfrak{a}_{\bullet}(v))}<+\infty.$
Then there exists a divisorial valuation $w\in\mathrm{Val}_{X,x}^{\mathrm{div}}$ such that $w(\mathfrak{a}_{\bullet}(v))>0$. By rescaling, we may assume that $w(\mathfrak{a}_{\bullet}(v))=1$, so that $w(f)\geq v(f)$ for all $f\in\mathscr{O}_{X,x}$. Thus
$\mathrm{vol}_{X}(v)\geq\mathrm{vol}_{X}(w)>0.$
By [BLQ_filt, Prop. 3.12], $v(\tilde{\mathfrak{a}}_{\bullet})=v(\mathfrak{a}_{\bullet})$. Hence $v$ also computes the lc threshold of $\tilde{\mathfrak{a}}_{\bullet}$. ∎
2.3.2. Normalized volumes and the minimizers
Definition 2.16 ([Li_vol]).
Let $x\in(X,\Delta)$ be a klt singularity, and $v\in\mathrm{Val}_{X,x}$ be a real valuation. The normalized volume of $v$ is
$\widehat{\mathrm{vol}}(v)=\widehat{\mathrm{vol}}_{X,\Delta}(v)\coloneqq A_{X,\Delta}(v)^{n}\cdot\mathrm{vol}_{X}(v)\in\mathbb{R}_{\geq 0}\cup\{+\infty\}$
where $n=\dim_{x}(X)$, and by convention $+\infty\cdot 0=+\infty$.
The local volume of $x\in(X,\Delta)$ is
$\widehat{\mathrm{vol}}(x;X,\Delta)\coloneqq\inf\mathopen{}\mathclose{{}\left\{\widehat{\mathrm{vol}}_{X,\Delta}(v):v\in\mathrm{Val}_{X,x}}\right\}.$
The local volume is positive by [Li_vol, Thm. 1.2]. A real valuation $v\in\mathrm{Val}_{X,x}$ is called a minimizer for the normalized volume function if $\widehat{\mathrm{vol}}(v)=\widehat{\mathrm{vol}}(x;X,\Delta)$.
Let $x\in(X,\Delta)$ be a klt singularity. Then there is a minimizer $v\in\mathrm{Val}_{X,x}$ of the normalized function by [Blu_exist, Main Thm.] and [Xu_qm, Rmk. 3.8], which is unique up to scaling by [XZ_unique, Thm. 1.1] and [BLQ_filt, Cor. 1.3], and is quasi-monomial by [Xu_qm, Thm. 1.2].
Lemma 2.17.
Let $x\in(X,\Delta)$ be a klt singularity, and $v^{\mathrm{m}}\in\mathrm{Val}_{X,x}^{\mathrm{qm}}$ be a minimizer of the normalized volume function. Let $n=\dim_{x}(X)$. Then the following hold:
  1. (1)
    $v^{\mathrm{m}}$ computes the lc threshold of $\mathfrak{a}_{\bullet}(v^{\mathrm{m}})$, and
    $\widehat{\mathrm{vol}}(x;X,\Delta)=\mathrm{lct}(X,\Delta;\mathfrak{a}_{\bullet}(v^{\mathrm{m}}))^{n}\cdot\mathrm{e}_{X}(\mathfrak{a}_{\bullet}(v^{\mathrm{m}}))=\inf\mathopen{}\mathclose{{}\left\{\mathrm{lct}(X,\Delta;\mathfrak{a}_{\bullet})^{n}\cdot\mathrm{e}_{X}(\mathfrak{a}_{\bullet})}\right\}$
    where $\mathfrak{a}_{\bullet}$ ranges over all ideal sequences centered at $x$;
  2. (2)
    if $v\in\mathrm{Val}_{X,x}$ computes the lc threshold of $\mathfrak{a}_{\bullet}(v^{\mathrm{m}})$, then $v=c\cdot v^{\mathrm{m}}$ for some $c\in\mathbb{R}_{>0}$;
  3. (3)
    if $\mathfrak{a}_{\bullet}=\{\mathfrak{a}_{\lambda}\}_{\lambda}$ is an ideal sequence centered at $x$ such that $\widehat{\mathrm{vol}}(x;X,\Delta)=\mathrm{lct}(X,\Delta;\mathfrak{a}_{\bullet})^{n}\cdot\mathrm{e}_{X}(\mathfrak{a}_{\bullet})$, then $\tilde{\mathfrak{a}}_{\lambda}=\mathfrak{a}_{c\lambda}(v^{\mathrm{m}})$ for some $c\in\mathbb{R}_{>0}$, for all $\lambda\geq 0$.
Proof.
(1). By [Liu_vol_KE, Thm. 27], we have
$\widehat{\mathrm{vol}}(x;X,\Delta)=\inf\mathopen{}\mathclose{{}\left\{\mathrm{lct}(X,\Delta;\mathfrak{a}_{\bullet})^{n}\cdot\mathrm{e}_{X}(\mathfrak{a}_{\bullet})}\right\}.$
Since $v^{\mathrm{m}}(\mathfrak{a}_{\bullet}(v^{\mathrm{m}}))=1$, and $\mathrm{e}_{X}(\mathfrak{a}_{\bullet}(v^{\mathrm{m}}))=\mathrm{vol}_{X}(v^{\mathrm{m}})$ by definition, and
$\mathrm{lct}(X,\Delta;\mathfrak{a}_{\bullet}(v^{\mathrm{m}}))\leq\frac{A_{X,\Delta}(v^{\mathrm{m}})}{v^{\mathrm{m}}(\mathfrak{a}_{\bullet}(v^{\mathrm{m}}))}=A_{X,\Delta}(v^{\mathrm{m}}),$
we have
$\mathrm{lct}(X,\Delta;\mathfrak{a}_{\bullet}(v^{\mathrm{m}}))^{n}\cdot\mathrm{e}_{X}(\mathfrak{a}_{\bullet}(v^{\mathrm{m}}))\leq A_{X,\Delta}(v^{\mathrm{m}})^{n}\cdot\mathrm{vol}_{X}(v^{\mathrm{m}})=\widehat{\mathrm{vol}}(x;X,\Delta).$
Thus the equality holds, and $v^{\mathrm{m}}$ computes the lc threshold $\mathrm{lct}(X,\Delta;\mathfrak{a}_{\bullet}(v^{\mathrm{m}}))$.
(2). Suppose $v$ computes the lc threshold of $\mathfrak{a}_{\bullet}(v^{\mathrm{m}})$, that is,
$\mathrm{lct}(X,\Delta;\mathfrak{a}_{\bullet}(v^{\mathrm{m}}))=\frac{A_{X,\Delta}(v)}{v(\mathfrak{a}_{\bullet}(v^{\mathrm{m}}))}.$
By scaling, we may assume that $v(\mathfrak{a}_{\bullet}(v^{\mathrm{m}}))=\inf v(\mathfrak{a}_{\lambda}(v^{\mathrm{m}}))/\lambda=1$. Thus $v(\mathfrak{a}_{\lambda}(v^{\mathrm{m}}))\geq\lambda$, i.e.,
$\mathfrak{a}_{\lambda}(v^{\mathrm{m}})\subseteq\mathfrak{a}_{\lambda}(v)$
for all $\lambda\geq 0$. Hence
$\mathrm{vol}_{X}(v)=\mathrm{e}_{X}(\mathfrak{a}_{\lambda}(v))\leq\mathrm{e}_{X}(\mathfrak{a}_{\lambda}(v^{\mathrm{m}}))=\mathrm{vol}_{X}(v^{\mathrm{m}}).$
Since $A_{X,\Delta}(v)=\mathrm{lct}(X,\Delta;\mathfrak{a}_{\bullet}(v^{\mathrm{m}}))=A_{X,\Delta}(v^{\mathrm{m}})$, we get
$\widehat{\mathrm{vol}}_{X,\Delta}(v)\leq\widehat{\mathrm{vol}}_{X,\Delta}(v^{\mathrm{m}})=\widehat{\mathrm{vol}}(x;X,\Delta).$
By uniqueness of the minimizer, $v=c\cdot v^{\mathrm{m}}$ for some $c>0$. In fact, $v=v^{\mathrm{m}}$ since $v(\mathfrak{a}_{\bullet}(v^{\mathrm{m}}))=1$.
(3). This is obtained in the proof of [BLQ_filt, Cor. 1.3], which we briefly recall: We may assume that $\mathrm{lct}(\mathfrak{a}_{\bullet})=\mathrm{lct}(\mathfrak{a}_{\bullet}(v^{\mathrm{m}}))=1$ by scaling. Let $\mathfrak{a}_{\bullet,t}$, where $0\leq t\leq 1$, be the geodesic between $\mathfrak{a}_{\bullet,0}=\mathfrak{a}_{\bullet}$ and $\mathfrak{a}_{\bullet,1}=\mathfrak{a}_{\bullet}(v^{\mathrm{m}})$ defined in [BLQ_filt, Def. 4.1]. Then
$\mathrm{lct}(X,\Delta;\mathfrak{a}_{\bullet,t})\leq(1-t)\cdot\mathrm{lct}(\mathfrak{a}_{\bullet})+t\cdot\mathrm{lct}(\mathfrak{a}_{\bullet}(v^{\mathrm{m}}))=1$
by [BLQ_filt, Prop. 5.1], and
$\mathrm{e}_{X}(\mathfrak{a}_{\bullet,t})^{-1/n}\geq(1-t)\cdot\mathrm{e}_{X}(\mathfrak{a}_{\bullet})^{-1/n}+t\cdot\mathrm{e}_{X}(\mathfrak{a}_{\bullet}(v^{\mathrm{m}}))^{-1/n}=\widehat{\mathrm{vol}}(x;X,\Delta)^{-1/n}$
by [BLQ_filt, Thm. 1.1(2)]. Hence
$\mathrm{lct}(X,\Delta;\mathfrak{a}_{\bullet,t})^{n}\cdot\mathrm{e}_{X}(\mathfrak{a}_{\bullet,t})\leq\widehat{\mathrm{vol}}(x;X,\Delta),$
so all inequalities above are equalities. Then $\mathfrak{a}_{\bullet}$ and $\mathfrak{a}_{\bullet}(v^{\mathrm{m}})$ have the same saturation, up to scaling, by [BLQ_filt, Thm. 1.1(3)]. But $\mathfrak{a}_{\bullet}(v^{\mathrm{m}})$ is saturated by [BLQ_filt, Lem. 3.20], so we get (3). ∎
3. Special valuations
3.1. Models of qdlt Fano type
Definition 3.1 ([XZ_stable_deg, Def. 3.5, 3.7]).
Let $(X,\Delta)$ be an affine klt pair. A model
$f\colon(Y,E=E_{1}+\cdots+E_{r})\to(X,\Delta)$
is said to be of qdlt Fano type if there is an effective $\mathbb{Q}$-divisor $D$ on $Y$ such that $\lfloor D\rfloor=0$, $D\geq f^{-1}_{*}\Delta$, $(Y,D+E)$ is qdlt, and $-(K_{Y}+D+E)$ is $f$-ample.
The model $f\colon(Y,E)\to(X,\Delta)$ is called a qdlt anti-canonical model if one can take $D=f^{-1}_{*}\Delta$ above, that is, $(Y,f^{-1}_{*}\Delta+E)$ is qdlt and $-(K_{Y}+f^{-1}_{*}\Delta+E)$ is $f$-ample.
Let $x\in(X,\Delta)$ be a klt singularity. A qdlt anti-canonical model $f\colon(Y,E)\to(X,\Delta)$ is called a Kollár model at $x$ if all components of $E$ are centered at $x$.
Remark.
If $f\colon(Y,E)\to(X,\Delta)$ is a Kollár model at $x\in(X,\Delta)$ and $E$ is irreducible, then $E$ is a Kollár component of $x\in(X,\Delta)$, and $f$ is the corresponding plt blow-up.
Lemma 3.2 ([XZ_stable_deg, Prop. 3.9]).
Let $(X,\Delta)$ be an affine klt pair, and $f\colon(Y,E)\to(X,\Delta)$ be a model of qdlt Fano type. Then there exists a birational contraction $\phi\colon Y\dashrightarrow\overline{Y}$ over $X$ that is a local isomorphism at generic points of strata of $(Y,E)$, such that $\bar{f}\colon(\overline{Y},\overline{E}=\phi_{*}E)\to(X,\Delta)$ is a qdlt anti-canonical model.
Lemma 3.3.
Let $(X,\Delta)$ be an affine klt pair, and $\bar{f}\colon(\overline{Y},\overline{E}=\sum_{i=1}^{r}\overline{E}_{i})\to(X,\Delta)$ be a qdlt anti-canonical model. Suppose $f\colon(Y,E)\to(X,\Delta)$ is a model such that $E=\sum_{i=1}^{r}E_{i}$ contains all $f$-exceptional divisors, and $\mathrm{ord}_{E_{i}}=\mathrm{ord}_{\overline{E}_{i}}$ for each $i$. Then $(Y,E)$ is of qdlt Fano type, and the birational map $\phi\colon Y\dashrightarrow\overline{Y}$ over $X$ is a local isomorphism at each stratum of $(Y,E)$.
Proof.
Note that the birational map $\phi=\bar{f}^{-1}\circ f\colon Y\dashrightarrow\overline{Y}$ is an isomorphism in codimension $1$. Thus if $-(K_{Y}+f^{-1}_{*}\Delta+E)$ is $f$-ample, then $\phi$ is an isomorphism.
Assume that $-(K_{Y}+f^{-1}_{*}\Delta+E)$ is $f$-semi-ample, then $\phi\colon Y\to\overline{Y}$ is the ample model for it over $X$, so in particular $\phi$ is a morphism. Since $\phi$ has no exceptional divisors, it is a local isomorphism over the locus where $\overline{Y}$ is $\mathbb{Q}$-factorial, which contains all generic points of strata of $(\overline{Y},\overline{E})$ as $(\overline{Y},\bar{f}^{-1}_{*}\Delta+\overline{E})$ is qdlt. Thus $f\colon(Y,E)\to(X,\Delta)$ is of qdlt Fano type by [XZ_stable_deg, Prop. 3.6].
In general, choose a general $\mathbb{Q}$-divisor $\overline{G}\in|{-(K_{\overline{Y}}+\bar{f}^{-1}_{*}\Delta+\overline{E})}|_{\mathbb{Q}}$ such that $(\overline{Y},\bar{f}^{-1}_{*}\Delta+\overline{G}+\overline{E})$ is qdlt and $\lfloor\bar{f}^{-1}_{*}\Delta+\overline{G}+\overline{E}\rfloor=\overline{E}$. Let $\Gamma=\bar{f}_{*}\overline{G}$, then $(X,\Delta+\Gamma)$ is lc,
$\bar{f}^{*}(K_{X}+\Delta+\Gamma)=K_{\overline{Y}}+\bar{f}^{-1}_{*}\Delta+\overline{G}+\overline{E},$
and $\mathrm{LCP}(X,\Delta+\Gamma)=\mathrm{QM}(\overline{Y},\overline{E})$. In particular, $A_{X,\Delta+\Gamma}(E_{i})=A_{X,\Delta+\Gamma}(\overline{E}_{i})=0$ for all $i$ since $E_{i}$ and $\overline{E}_{i}$ corresponds to the same divisorial valuation on $X$. Hence we have
$f^{*}(K_{X}+\Delta+(1-t)\Gamma)=K_{Y}+f^{-1}_{*}(\Delta+(1-t)\Gamma)+\sum_{i=1}^{r}(1-t\mathop{\mathrm{ord}_{E_{i}}}(\Gamma))E_{i}$
for all $t\geq 0$. Thus $(Y,f^{-1}_{*}(\Delta+\Gamma)+E)$ is lc, and $Y$ is of Fano type over $X$.
Since $(Y,f^{-1}_{*}\Delta+E)$ is potentially lc, by [Kol_log-plurigenera, Prop. 19], there exists a small lc modification
$\mu\colon(Y^{\mathrm{c}},(f^{\mathrm{c}})^{-1}_{*}\Delta+E^{\mathrm{c}})\to(Y,f^{-1}_{*}\Delta+E),$
where $f^{\mathrm{c}}=f\circ\mu$, $E^{\mathrm{c}}=\mu^{-1}_{*}E$, and $\mu$ is a local isomorphism at every lc center of $(Y^{\mathrm{c}},(f^{\mathrm{c}})^{-1}_{*}\Delta+E^{\mathrm{c}})$. Then it suffices to show that $f^{\mathrm{c}}\colon(Y^{\mathrm{c}},E^{\mathrm{c}})\to(X,\Delta)$ is of qdlt Fano type by [XZ_stable_deg, Prop. 3.6].
Hence we may assume that $-(K_{Y}+f^{-1}_{*}\Delta+E)$ is $\mathbb{Q}$-Cartier. Since $Y$ is of Fano type over $X$, we can run a $-(K_{Y}+f^{-1}_{*}\Delta+E)$-MMP and get a good minimal model
[…]
over $X$. Note that $\bar{f}\colon\overline{Y}\to X$ is the ample model for $-(K_{Y}+f^{-1}_{*}\Delta+E)$ over $X$, so it is also the ample for $-(K_{Y^{\prime}}+f^{\prime-1}_{*}\Delta+\psi_{*}E)$. Since $\phi$ is small, none of the component $E_{i}$ is contracted by $\psi$. Now by the relatively semi-ample case we proved above, $f^{\prime}\colon(Y^{\prime},E^{\prime}=\psi_{*}E)\to(X,\Delta)$ is a model of qdlt Fano type. In particular, $(Y^{\prime},f^{\prime-1}_{*}\Delta+E^{\prime})$ is qdlt. By the MMP we have
$A_{Y^{\prime},f^{\prime-1}_{*}\Delta+E^{\prime}}(F)\leq A_{Y,f^{-1}_{*}\Delta+E}(F)$
for every prime divisor $F$, and the strict inequality holds if and only if $\psi$ is not a local isomorphism at $\mathrm{center}_{Y}(F)$. Hence $\psi$ is a local isomorphism at every lc center $z\in(Y,f^{-1}_{*}\Delta+E)$, and $\psi(z)$ is also an lc center of $(Y^{\prime},f^{\prime-1}_{*}\Delta+E)$. Thus $(Y,f^{-1}_{*}\Delta+E)$ is simple-toroidal at $z$. So $(Y,f^{-1}_{*}\Delta+E)$ is qdlt.
Since $f^{\prime}\colon(Y^{\prime},E^{\prime})\to(X,\Delta)$ is of qdlt Fano type, by definition there exists a $\mathbb{Q}$-divisor $D^{\prime}\geq f^{\prime-1}_{*}\Delta$ on $Y^{\prime}$ such that $\lfloor D^{\prime}+E^{\prime}\rfloor=E^{\prime}$, $(Y^{\prime},D^{\prime}+E^{\prime})$ is qdlt, and $-(K_{Y^{\prime}}+D^{\prime}+E^{\prime})$ is $f^{\prime}$-ample. Let $A$ be an $f$-ample divisor on $Y$, and $A^{\prime}=\psi_{*}A$. Replacing $A$ with $\epsilon A$ for some $0<\epsilon\ll 1$, we may assume that $-(K_{Y^{\prime}}+D^{\prime}+E^{\prime})-A^{\prime}$ is $f^{\prime}$-ample. Let
$H^{\prime}\in\mathopen{}\mathclose{{}\left|{-(K_{Y^{\prime}}+D^{\prime}+E^{\prime})-A^{\prime}}}\right|_{\mathbb{Q}}$
be a general $\mathbb{Q}$-divisor, so that $H^{\prime}$ does not contain any stratum of $(Y^{\prime},E^{\prime})$. Thus $H=\psi^{-1}_{*}H^{\prime}$ does not contain any stratum of $(Y,E)$ since $\psi$ is a local isomorphism at the generic points thereof, and nor does $D=\psi^{-1}_{*}D^{\prime}$. Hence
$-(K_{Y}+D+H+E)\sim_{\mathbb{Q}}A$
is $f$-ample, and $(Y,f^{-1}_{*}\Delta+\epsilon(D+H-f^{-1}_{*}\Delta)+E)$ is qdlt for some $0<\epsilon\ll 1$. Recall that
$K_{Y}+f^{-1}_{*}(\Delta+\Gamma)+E\sim_{X,\mathbb{Q}}0$
and $(Y,f^{-1}_{*}(\Delta+\Gamma)+E)$ is lc. Let
$D_{Y}$ $=f^{-1}_{*}\Delta+\epsilon(D+H-f^{-1}_{*}\Delta)+(1-\epsilon)f^{-1}_{*}\Gamma$
$=\epsilon(D+H)+(1-\epsilon)f^{-1}_{*}(\Delta+\Gamma)$
then $D_{Y}\geq f^{-1}_{*}\Delta$, $\lfloor D_{Y}\rfloor=0$,
$-(K_{Y}+D_{Y}+E)\sim_{X,\mathbb{Q}}\epsilon A$
is $f$-ample, and $(Y,D_{Y}+E)$ is qdlt. That is, $f\colon(Y,E)\to(X,\Delta)$ is of qdlt Fano type. ∎
Definition 3.4 ([XZ_stable_deg, Def. 3.12]).
Let $S$ be a reduced scheme, and $\pi\colon x\in(X,\Delta)\to S$ be a locally stable family of klt singularities. A family of models $f\colon(Y,E)\to(X,\Delta)$ at $x$ is called a locally stable family of Kollár models at $x$ if $\pi\circ f\colon(Y,f^{-1}_{*}\Delta+E)\to S$ is a locally stable family with qdlt fibers, and $-(K_{Y}+f^{-1}_{*}\Delta+E)$ is $f$-ample.
Note that the fiber $f_{s}\colon(Y_{s},E_{s})\to(X_{s},\Delta_{s})$ is a Kollár model at $x_{s}$ for every $s\in S$.
Remark.
Let $S$ be a regular local scheme of dimension $d$, with the generic point $\eta\in S$ and the closed point $s\in S$. Let $H=\sum_{j=1}^{d}H_{j}$ be an snc divisor on $S$ defined by a regular system of parameters.
Suppose $(X,\Delta)$ is a pair, and $\pi\colon X\to S$ is a dominant morphism that is separated and of finite type. Then $\pi\colon(X,\Delta)\to S$ is a locally stable family if and only if $(X,\Delta+\pi^{*}H)$ is slc; see [Kol_fam, Thm. 4.54]. By inversion of adjunction, $\pi\colon(X,\Delta)\to S$ is a locally stable family of qdlt (resp., klt) pairs if and only if $(X,\Delta+\pi^{*}H)$ is qdlt (resp., qdlt with $\lfloor\Delta\rfloor=0$).
Suppose $\pi\colon x\in(X,\Delta)\to S$ is a locally stable family of klt singularities, then a locally stable family of Kollár models
$f\colon(Y,E)\to(X,\Delta)$
is a model for $(X,\Delta)$ such that each component of $E$ is centered on $X_{\eta}$, $(Y,f^{-1}_{*}\Delta+E+f^{*}\pi^{*}H)$ is qdlt, and $-(K_{Y}+f^{-1}_{*}\Delta+E)$ is $f$-ample. In other words,
$f\colon(Y,E+f^{*}\pi^{*}H)\to(X,\Delta)$
is a qdlt anti-canonical model for $(X,\Delta)$.
3.2. Special valuations and complements
Definition 3.5.
Let $(X,\Delta)$ be a pair. A subset $\sigma\subseteq\mathrm{Val}_{X}$ is called a quasi-monomial (simplicial) cone for $(X,\Delta)$ if $\sigma=\mathrm{QM}_{y}(Y,E)$, where $f\colon(Y,E)\to(X,\Delta)$ is a model, $E$ contains all $f$-exceptional divisors, and $y\in(Y,E)$ is a generic point of a stratum such that $(Y,f^{-1}_{*}\Delta+E)$ is simple-toroidal at $y$.
Note that by taking a log resolution $(Y^{\prime},E^{\prime})\to(Y,f^{-1}_{*}\Delta+E)$ that is a local isomorphism over $y$, we may assume that $(Y,E)$ is a simple-toroidal model.
If $E_{1},\ldots,E_{r}$ are the components of $E$ that contain $y$, then there is an isomorphism $\sigma\simeq\mathbb{R}_{\geq 0}^{r}$. If $\tau\subseteq\sigma$ corresponds to a rational simplicial cone in $\mathbb{R}_{\geq 0}^{r}$, then $\tau\subseteq\mathrm{Val}_{X}$ is also a quasi-monomial simplicial cone for $(X,\Delta)$. If $K_{X}+\Delta$ is $\mathbb{Q}$-Cartier, then the log discrepancy function $A_{X,\Delta}$ is linear on $\sigma$.
Definition 3.6.
Let $(X,\Delta)$ be an affine klt pair, and $f\colon(Y,E)\to(X,\Delta)$ be a simple-toroidal model. A $\mathbb{Q}$-complement $\Gamma$ of $(X,\Delta)$ is said to be special with respect to $f\colon(Y,E)\to(X,\Delta)$ if $f^{-1}_{*}\Gamma\geq G$ for some effective $f$-ample $\mathbb{Q}$-divisor $G$ on $Y$ whose support does not contain any stratum of $(Y,E)$.
A quasi-monomial simplicial cone $\sigma\subseteq\mathrm{Val}_{X}$ for $(X,\Delta)$ is said to be special if there is a simple-toroidal model $f\colon(Y,E)\to(X,\Delta)$, a special $\mathbb{Q}$-complement $\Gamma$ of $(X,\Delta)$ with respect to $f\colon(Y,E)\to(X,\Delta)$, and a generic point $y$ of a stratum of $(Y,E)$, such that $\sigma\subseteq\mathrm{QM}_{y}(Y,E)\cap\mathrm{LCP}(X,\Delta+\Gamma)$.
A real valuation $v\in\mathrm{Val}_{X}$ is said to be special for $(X,\Delta)$ if $v\in\sigma$ for some special quasi-monomial simplicial cone $\sigma\subseteq\mathrm{Val}_{X}$. In particular, $v$ is quasi-monomial.
Lemma 3.7.
Let $(X,\Delta)$ be an affine klt pair, and $\sigma\subseteq\mathrm{Val}_{X}$ be a quasi-monomial cone for $(X,\Delta)$. Then the following are equivalent:
  1. (i)
    $\sigma$ is special.
  2. (ii)
    there exists a model $f\colon(Y,E)\to(X,\Delta)$ of qdlt Fano type such that $\sigma=\mathrm{QM}(Y,E)$;
Proof.
(i) $\Rightarrow$ (ii). See [XZ_stable_deg, Thm. 3.14] or [Xu_K-Stability_Book, Thm. 5.26].
(ii) $\Rightarrow$ (i). By Lemma 3.2, we may assume $(Y,f^{-1}_{*}\Delta+E)$ is qdlt and $-(K_{Y}+f^{-1}_{*}\Delta+E)$ is ample. We can choose a general $\mathbb{Q}$-divisor $G\in|{-(K_{Y}+f^{-1}_{*}\Delta+E)}|_{\mathbb{Q}}$ such that $(Y,f^{-1}_{*}\Delta+G+E)$ is qdlt, and $\lfloor f^{-1}_{*}\Delta+G+E\rfloor=E$. Note that $f_{*}(G+E)\sim_{\mathbb{Q}}-(K_{X}+\Delta)$, where $f_{*}E$ consists of components of $E$ that are not $f$-exceptional. Let
$\mu\colon(Z,F)\to(Y,f^{-1}_{*}\Delta+G+E)$
be a log resolution such that $\mu$ is a local isomorphism over the simple-toroidal locus of $(Y,E)$, and there is a $\mu$-ample divisor $-A$ on $Z$ with $A\geq 0$ and $\mathrm{Supp}(A)=\mathrm{Ex}(\mu)$. Then we have
$0\sim_{\mathbb{Q}}\mu^{*}(K_{Y}+f^{-1}_{*}\Delta+G+E)=K_{Z}+\mu^{-1}_{*}f^{-1}_{*}\Delta+(\mu^{*}G-\epsilon A)+\mu^{-1}_{*}E+F^{\prime}+\epsilon A.$
where $F^{\prime}=\sum_{i}A_{Y,f^{-1}_{*}\Delta+E}(F_{i})\cdot F_{i}$ and $F_{i}$ ranges over all $\mu$-exceptional divisors. By our choice of $\mu$, we have $\lfloor F^{\prime}\rfloor\leq 0$. Thus $\lfloor F^{\prime}+\epsilon A\rfloor\leq 0$, and $\mu^{*}G-\epsilon A$ is ample for $0<\epsilon\ll 1$. Now
$\mathrm{LCP}(Z,\mu^{-1}_{*}f^{-1}_{*}\Delta+\mu^{-1}_{*}E+F^{\prime}+\epsilon A)=\mathrm{QM}(Z,\mu^{-1}_{*}E)=\mathrm{QM}(Y,E).$
Choose a general $H\in|\mu^{*}G-\epsilon A|_{\mathbb{Q}}$, such that $(Z,\mu^{-1}_{*}f^{-1}_{*}\Delta+H+\mu^{-1}_{*}E+F^{\prime}+\epsilon A)$ is sub-lc and
$\mathrm{LCP}(Z,\mu^{-1}_{*}f^{-1}_{*}\Delta+H+\mu^{-1}_{*}E+F^{\prime}+\epsilon A)=\mathrm{LCP}(Z,\mu^{-1}_{*}f^{-1}_{*}\Delta+\mu^{-1}_{*}E+F^{\prime}+\epsilon A).$
Let $\Gamma=f_{*}\mu_{*}H+f_{*}E\sim_{\mathbb{Q}}f_{*}(G+E)\sim_{\mathbb{Q}}-(K_{X}+\Delta)$, then
$\mu^{*}f^{*}(K_{X}+\Delta+\Gamma)=K_{Z}+\mu^{-1}_{*}f^{-1}_{*}\Delta+H+\mu^{-1}_{*}E+F^{\prime}+\epsilon A.$
Hence $\Gamma$ is a special $\mathbb{Q}$-complement with respect to $(Z,F)\to(X,\Delta)$, and $\mathrm{LCP}(X,\Delta+\Gamma)=\sigma$. ∎
Lemma 3.8.
Let $x\in(X,\Delta)$ be a klt singularity. Suppose $v\in\mathrm{Val}_{X,x}^{\mathrm{qm}}$, and let $\mathfrak{a}_{\bullet}=\mathfrak{a}_{\bullet}(v)$ denote the associated ideal sequence. Then the following are equivalent:
  1. (i)
    $v$ is special;
  2. (ii)
    there exists a Kollár model $f\colon(Y,E)\to(X,\Delta)$ at $x$ such that $v\in\mathrm{QM}(Y,E)$;
  3. (iii)
    the associated graded ring
    $R_{v}=\mathrm{gr}_{v}(\mathscr{O}_{X})=\bigoplus_{\lambda\geq 0}\mathfrak{a}_{\lambda}/\mathfrak{a}_{>\lambda}$
    is a finitely generated graded $\kappa(x)$-algebra, and the pair $(X_{v}=\mathop{\mathrm{Spec}(R_{v})},\Delta_{v})$ is klt, where $\Delta_{v}$ is defined by the divisorial part of the initial ideal of $\Delta$.
  4. (iv)
    there exists $\delta>0$ such that if $D$ is an effective $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisor on $X$ with $v(D)<\delta$, then $(X,\Delta+D)$ is klt and $v$ is special for $(X,\Delta+D)$.
  5. (v)
    there exists $\delta>0$ such that if $D$ is an effective $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisor on $X$ with $v(D)<\delta$, then $(X,\Delta+D)$ is klt and $v$ computes the lc threshold $\mathrm{lct}(X,\Delta+D;\mathfrak{a}_{\bullet})$.
  6. (vi)
    for every effective $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisor $D$ on $X$, there exists a $\mathbb{Q}$-complement $\Gamma$ of $(X,\Delta)$ such that $\Gamma\geq\epsilon D$ for some $\epsilon>0$, and $v\in\mathrm{LCP}(X,\Delta+\Gamma)$.
Proof.
(i), (ii), and (iii) are equivalent by [XZ_stable_deg, Thm. 4.1].
(i-iii) $\Rightarrow$ (iv). The valuation $v$ induces a canonical valuation $\mathrm{wt}_{v}$ on $X_{v}$, such that $\mathrm{wt}_{v}(D_{v})=v(D)$ for every effective $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisor $D$ on $X$. By the Izumi type inequality [Zhuang_boundednessI, Lem. 3.5],
$\mathrm{lct}(X_{v},\Delta_{v};D_{v})\geq c_{0}\frac{\widehat{\mathrm{vol}}(x_{v};X_{v},\Delta_{v})}{\widehat{\mathrm{vol}}_{X_{v},\Delta_{v}}(\mathrm{wt}_{v})}\cdot\frac{A_{X_{v},\Delta_{v}}(\mathrm{wt}_{v})}{\mathrm{wt}_{v}(D_{v})},$
where $x_{v}\in X_{v}$ is the vertex defined by the irrelevant ideal of $R_{v}$, and $c_{0}$ is a constant only depending on $\dim_{x}(X)=\dim(X_{v})$. Thus, if
$v(D)=\mathrm{wt}_{v}(D_{v})<\delta\coloneqq c_{0}\frac{\widehat{\mathrm{vol}}(x_{v};X_{v},\Delta_{v})}{\widehat{\mathrm{vol}}_{X_{v},\Delta_{v}}(\mathrm{wt}_{v})}A_{X_{v},\Delta_{v}}(\mathrm{wt}_{v}),$
then $(X_{v},\Delta_{v}+D_{v})$ is klt. Since we know (iii) $\Rightarrow$ (i), this implies that $v$ is special for $(X,\Delta+D)$.
(iv) $\Rightarrow$ (v). It suffices to show that if $v$ is special for $(X,\Delta)$, then $v$ computes the lc threshold of $\mathfrak{a}_{\bullet}$. Suppose $v$ is special, then $v\in\mathrm{LCP}(X,\Delta+\Gamma)$ for some $\mathbb{Q}$-complement $\Gamma$. Hence we get the conclusion by [JM, Thm. 7.8]. We also give a proof here: For $w\in\mathrm{Val}_{X,x}$, we have
$A_{X,\Delta}(w)-w(\Gamma)=A_{X,\Delta+\Gamma}(w)\geq 0$
since $(X,\Delta+\Gamma)$ is lc, and $w(\mathfrak{a}_{\bullet})\leq w(\Gamma)/v(\Gamma)$. Meanwhile, $A_{X,\Delta}(v)=v(\Gamma)$ since $v\in\mathrm{LCP}(X,\Delta+\Gamma)$. Thus
$\frac{A_{X,\Delta}(w)}{w(\mathfrak{a}_{\bullet})}\geq\frac{A_{X,\Delta}(w)}{w(\Gamma)}v(\Gamma)\geq v(\Gamma)=A_{X,\Delta}(v)=\frac{A_{X,\Delta}(v)}{v(\mathfrak{a}_{\bullet})}.$
Taking infimum over all $w\in\mathrm{Val}_{X,x}$, we get $\mathrm{lct}(X,\Delta;\mathfrak{a}_{\bullet})=A_{X,\Delta}(v)/v(\mathfrak{a}_{\bullet})$.
(v) $\Rightarrow$ (vi). Let $\epsilon<\delta/v(D)$, so $(X,\Delta+\epsilon D)$ is klt and $v$ computes the lc threshold $\mathrm{lct}(X,\Delta+\epsilon D;\mathfrak{a}_{\bullet})$ by (v). Then the conclusion follows from the fact that a quasi-monomial valuation that computes the lc threshold of an ideal sequence is an lc place of a $\mathbb{Q}$-complement; see [LX_stability, Rmk. 4.4]. This fact is also the special case of Lemma 4.2 when the base $S$ is 00-dimensional.
(vi) $\Rightarrow$ (i). The proof is the same as [XZ_stable_deg, Lem. 3.4]. Let $f\colon(Y,E)\to(X,\Delta)$ be a simple-toroidal model, and $y\in(Y,E)$ be a generic point of a stratum, such that $v\in\mathrm{QM}_{y}(Y,E)$, and there is an $f$-ample divisor $-A$ with $A\geq 0$ and $\mathrm{Supp}(A)=\mathrm{Ex}(f)$. Then $-A$ is ample since $X$ is affine. Let $H\in|{-A}|_{\mathbb{Q}}$ be a general $\mathbb{Q}$-divisor whose support does not contain any stratum of $(Y,\mathrm{Supp}(f^{-1}_{*}\Delta+E+\mathrm{Ex}(f)))$. Let $D=f_{*}H$, then there is a $\mathbb{Q}$-complement $\Gamma\geq\epsilon D$ of $(X,\Delta)$ such that $v\in\mathrm{LCP}(X,\Delta+\Gamma)$ by (vi). By construction, $f^{-1}_{*}\Gamma\geq\epsilon H$. That is, $\Gamma$ is a special $\mathbb{Q}$-complement with respect to $f\colon(Y,E)\to(X,\Delta)$. ∎
Remark.
We have $A_{X_{v},\Delta_{v}}(\mathrm{wt}_{v})=A_{X,\Delta}(v)$ and $\widehat{\mathrm{vol}}_{X_{v},\Delta_{v}}(\mathrm{wt}_{v})=\widehat{\mathrm{vol}}_{X,\Delta}(v)$ by [LX_stability, Lem. 4.10]. Thus we can take
$\delta=c_{0}\frac{\widehat{\mathrm{vol}}(x_{v};X_{v},\Delta_{v})}{\widehat{\mathrm{vol}}_{X,\Delta}(v)}A_{X,\Delta}(v).$
in (iv) and (v). Moreover, if $v$ is the minimizer of the normalized volume function for $x\in(X,\Delta)$, then it is proved that one can take $\delta=A_{X,\Delta}(v)/(\dim X)$ in [XZ_stable_deg, Lem. 3.3] using K-stability of valuations, hence $v$ is special.
Lemma 3.9.
Let $x\in(X,\Delta)$ be a klt singularity. Suppose $v\in\mathrm{Val}_{X,x}$ is a special valuation for $(X,\Delta)$, and $\mathfrak{a}_{\bullet}=\mathfrak{a}_{\bullet}(v)$ is the ideal sequence associated to $v$. Then, up to scaling, $v$ is the unique valuation that computes the lc threshold $\mathrm{lct}(X,\Delta;\mathfrak{a}_{\bullet})$.
Proof.
We know $v$ computes the lc threshold of $\mathfrak{a}_{\bullet}$ by Lemma 3.8. Suppose $w\in\mathrm{Val}_{X,x}$ such that
$\frac{A_{X,\Delta}(v)}{v(\mathfrak{a}_{\bullet})}=\mathrm{lct}(X,\Delta;\mathfrak{a}_{\bullet})=\frac{A_{X,\Delta}(w)}{w(\mathfrak{a}_{\bullet})}.$
By rescaling, we may assume that $A_{X,\Delta}(v)=A_{X,\Delta}(w)=1$. Hence
$\inf_{\lambda>0}\frac{w(\mathfrak{a}_{\lambda}(v))}{\lambda}=w(\mathfrak{a}_{\bullet})=v(\mathfrak{a}_{\bullet})=1.$
Thus $w(f)\geq v(f)$ for every $f\in\mathscr{O}_{X}$. Assume that $w(f)>v(f)$ for some $f$, then $f\in\mathfrak{m}_{x}\smallsetminus\{0\}$, where $\mathfrak{m}_{x}\subseteq\mathscr{O}_{X}$ is the maximal ideal of $x$. Let $D=\mathrm{div}(f)$. Then, for $\epsilon>0$,
$\frac{A_{X,\Delta+\epsilon D}(w)}{w(\mathfrak{a}_{\bullet})}=1-\epsilon w(f)<1-\epsilon v(f)=\frac{A_{X,\Delta+\epsilon D}(v)}{v(\mathfrak{a}_{\bullet})}.$
This contradicts to that $v$ computes $\mathrm{lct}(X,\Delta+\epsilon D;\mathfrak{a}_{\bullet})$ for $0<\epsilon\ll 1$ by Lemma 3.8. ∎
Remark.
If $v=\mathrm{ord}_{E}$ is divisorial, and $f\colon(Y,E)\to(X,\Delta)$ is the prime blow-up where $-E$ is $f$-ample. Then the lemma follows from that $(Y,f^{-1}_{*}\Delta+E)$ is plt.
4. Families of Kollár models
4.1. Degeneration of valuations via families of models
In this subsection, let $S$ be a regular local scheme of dimension $d$, with the closed point $s\in S$ and the generic point $\eta\in S$. Let $H=\sum_{j=1}^{d}H_{j}$ be an snc divisor on $S$ defined by a regular system of parameters. The subscript $(-)_{t}$ will denote the fiber at a point $t\in S$ for objects over $S$.
Lemma 4.1.
Let $\pi\colon(X,\Delta)\to S$ be a locally stable family of pairs. Then there exists $\epsilon>0$ such that if $f\colon(Y,E)\to(X,\Delta)$ is a family of models where $E$ is $\mathbb{Q}$-Cartier and $(Y,f^{-1}_{*}\Delta+(1-\epsilon)E)\to S$ is locally stable, then $(Y,f^{-1}_{*}\Delta+E)\to S$ is also locally stable.
Proof.
This follows from the ACC for lc thresholds; see [HMX_ACC]. ∎
Lemma 4.2.
Let $\pi\colon x\in(X,\Delta)\to S$ be a locally stable family of klt singularities, and $I_{\bullet}\subseteq\mathscr{O}_{X}$ be an ideal sequence on $X$ cosupported on $\{x\}$ such that
$\mathrm{lct}(X_{s},\Delta_{s};I_{s,\bullet})=\mathrm{lct}(X_{\eta},\Delta_{\eta};I_{\eta,\bullet})<\infty$
Suppose $v_{0}\in\mathrm{Val}_{X_{\eta},x_{\eta}}^{\mathrm{qm}}$ computes the lc threshold of $I_{\eta,\bullet}$, and $\sigma\subseteq\mathrm{Val}_{X_{\eta},x_{\eta}}$ is a quasi-monomial simplicial cone for $(X_{\eta},\Delta_{\eta})$ such that $v_{0}\in\sigma$. Then there exists a family of models at $x$ over $S$
$f\colon(Y,E=E_{1}+\cdots+E_{r})\to(X,\Delta)$
such that the following hold:
  1. (1)
    $E_{1},\ldots,E_{r}$ are all the exceptional divisors of $f$;
  2. (2)
    the valuations $\mathrm{ord}_{E_{1,\eta}},\ldots,\mathrm{ord}_{E_{r,\eta}}$ span a simplicial cone $\tau\subseteq\sigma$ such that $v_{0}\in\tau$;
  3. (3)
    $\pi\circ f\colon(Y,f^{-1}_{*}\Delta+E)\to S$ is a locally stable family.
  4. (4)
    there exists a relative $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisor $\Gamma$ on $X/S$ such that $\pi\colon(X,\Delta+\Gamma)\to S$ is a locally stable family, $K_{X}+\Delta+\Gamma\sim_{S,\mathbb{Q}}0$, and $\tau\subseteq\mathrm{LCP}(X_{\eta},\Delta_{\eta}+\Gamma_{\eta})$.
Furthermore, we may assume that either $Y$ is $\mathbb{Q}$-factorial and $-(K_{Y}+f^{-1}_{*}\Delta+E)$ is $f$-semi-ample, or each $E_{i}$ is $\mathbb{Q}$-Cartier and there exists an $f$-ample divisor $-A$ on $Y$ with $A\geq 0$ and $\mathrm{Supp}(A)=E$.
Proof.
By rescaling the index of $I_{\bullet}$, we may assume that $\mathrm{lct}(X_{s},\Delta_{s};I_{s,\bullet})=\mathrm{lct}(X_{\eta},\Delta_{\eta};I_{\eta,\bullet})=1$. The assumption that $v$ computes the lc threshold means
$\mathrm{lct}(X_{\eta},\Delta_{\eta};I_{\eta,\bullet})=\frac{A_{X_{\eta},\Delta_{\eta}}(v_{0})}{v_{0}(I_{\eta,\bullet})}.$
By rescaling $v_{0}$, we may assume $A_{X_{\eta},\Delta_{\eta}}(v_{0})=1$. Hence $v_{0}(I_{\eta,\bullet})=1$. Since $X_{\eta}\subseteq X$ is the generic fiber, we identify $\mathrm{Val}_{X_{\eta},x_{\eta}}=\mathrm{Val}_{X,x}$, so that $A_{X,\Delta}(v_{0})=v_{0}(I_{\bullet})=1$. By inversion of adjunction, $(X,\Delta+\pi^{*}H)$ is dlt, and
$\mathrm{lct}(X,\Delta+\pi^{*}H;I_{\bullet})=\mathrm{lct}(X_{s},\Delta_{s};I_{s,\bullet})=1.$
It follows that $\mathrm{lct}(X,\Delta;I_{\bullet})=1$, and $v_{0}$ computes the lc threshold of $I_{\bullet}$ on $(X,\Delta)$.
Fix an isomorphism $\sigma\simeq\mathbb{R}_{\geq 0}^{r}$. Note that the function
$A_{X,\Delta+I_{\bullet}}(v)\coloneqq A_{X,\Delta}(v)-v(I_{\bullet})=A_{X,\Delta}(v)-\inf_{\lambda>0}\frac{v(I_{\lambda})}{\lambda}$
[…]
$A_{X,\Delta+I_{\bullet}}(v)\leq A_{X,\Delta+I_{\bullet}}(v_{0})+C\cdot\mathrm{d}(v_{0},v)=C\cdot\mathrm{d}(v_{0},v).$
Fix $10<\epsilon<$ 1 as in Lemma 4.1 for $\pi\colon(X,\Delta)\to S$. By Diophantine approximation (see [LX_stability, Lem. 2.7] or [Xu_K-Stability_Book, Lem. 4.47]), there exist $v_{1},\ldots,v_{r}\in\sigma$ and $q_{1},\ldots,q_{r}\in\mathbb{Z}_{>0}$ such that:
  1. (1)
    $v_{0}$ is in the convex cone $\tau$ spanned by $v_{1},\ldots,v_{r}$;
  2. (2)
    $q_{i}v_{i}=c_{i}\mathop{\mathrm{ord}_{E_{i}}}$, where $E_{i}$ is a prime divisor over $X$ and $c_{i}\in\mathbb{Z}_{>0}$, for all $i=1,\ldots,r$;
  3. (3)
    $\mathrm{d}(v_{0},v_{i})<\frac{\epsilon}{2Cq_{i}}$ for all $i=1,\ldots,r$.
Thus
$0\leq A_{X,\Delta+I_{\bullet}}(E_{i})=\frac{q_{i}}{c_{i}}A_{X,\Delta+I_{\bullet}}(v_{i})<\frac{\epsilon}{2}.$
for all $i=1,\ldots,r$.
Choose $\epsilon^{\prime}>0$ such that $2\epsilon^{\prime}\mathop{\mathrm{ord}_{E_{i}}}(I_{\bullet})<\epsilon$ for all $i=1,\ldots,r$. Since
$\sup_{\lambda>0}\lambda\cdot\mathrm{lct}(X_{s},\Delta_{s};I_{s,\lambda})=\lim_{\lambda\to\infty}\lambda\cdot\mathrm{lct}(X_{s},\Delta_{s};I_{s,\lambda})=\mathrm{lct}(X_{s},\Delta_{s};I_{s,\bullet})=1,$
we have
$(1-\epsilon^{\prime})/\lambda<\mathrm{lct}(X_{s},\Delta_{s};I_{s,\lambda})\leq 1/\lambda$
for $\lambda\gg 0$. Write $c=(1-\epsilon^{\prime})/\lambda$, then $\mathrm{lct}(X,\Delta;I_{\lambda}^{c})\geq\mathrm{lct}(X_{s},\Delta_{s};I_{s,\lambda}^{c})>1$, and
$A_{X,\Delta+I_{\lambda}^{c}}(E_{i})$ $\leq A_{X,\Delta+I_{\bullet}}(E_{i})+(\mathop{\mathrm{ord}_{E_{i}}}(I_{\bullet})-c\mathop{\mathrm{ord}_{E_{i}}}(I_{\lambda}))$
$<\frac{\epsilon}{2}+(1-\epsilon^{\prime})\mathopen{}\mathclose{{}\left(\mathop{\mathrm{ord}_{E_{i}}}(I_{\bullet})-\frac{1}{\lambda}\mathop{\mathrm{ord}_{E_{i}}}(I_{\lambda})}\right)+\epsilon^{\prime}\mathop{\mathrm{ord}_{E_{i}}}(I_{\bullet})$
$<\frac{\epsilon}{2}+0+\frac{\epsilon}{2}=\epsilon.$
where we used the general fact that $v(I_{\bullet})\leq v(I_{\lambda})/\lambda$ for every valuation $v$.
Write $a_{i}\coloneqq A_{X,\Delta+I_{\lambda}^{c}}(E_{i})<\epsilon<1$. By Lemma 2.4, there exists a model
$f\colon(Y,E)\to(X,\Delta)$
such that $E=E_{1}+\cdots+E_{r}$ is the sum of all exceptional divisors of $f$, and we may assume that either $Y$ is $\mathbb{Q}$-factorial, or each $E_{i}$ is $\mathbb{Q}$-Cartier and there exists an $f$-ample divisor $-A$ on $Y$ with $A\geq 0$ and $\mathrm{Supp}(A)=E$. Since
$\mathrm{lct}(X,\Delta+\pi^{*}H;I_{\lambda}^{c})=\mathrm{lct}(X_{s},\Delta_{s};I_{s,\lambda}^{c})>1,$
$(Y,f^{-1}_{*}\Delta+\sum_{i=1}^{r}(1-a_{i})E_{i}+f^{*}\pi^{*}H)$ is lc, that is, $\pi\circ f\colon(Y,f^{-1}_{*}\Delta+\sum_{i=1}^{r}(1-a_{i})E_{i})\to S$ is locally stable. Thus $\pi\circ f\colon(Y,f^{-1}_{*}\Delta+E)\to S$ is locally stable since $a_{i}<\epsilon$ for all $i$, by Lemma 4.1.
Now we get (1), (2), and (3). Choose a model $f\colon(Y,E)\to(X,\Delta)$ as above such that $Y$ is $\mathbb{Q}$-factorial. Since $Y$ is of Fano type over $X$, so we can run a $-(K_{Y}+f^{-1}_{*}\Delta+E)$-MMP over $X$ and get a $\mathbb{Q}$-factorial good minimal model
[…]
where $-(K_{Y^{\prime}}+f^{\prime-1}_{*}\Delta+\phi_{*}E)$ is $f$-semi-ample. As above, $(Y^{\prime},f^{\prime-1}_{*}\Delta+\sum_{i=1}^{r}(1-a_{i})\phi_{*}E_{i}+f^{\prime*}\pi^{*}H)$ is also lc, hence $\pi\circ f^{\prime}\colon(Y^{\prime},f^{\prime-1}_{*}\Delta+\phi_{*}E)\to S$ is locally stable by Lemma 4.1. In the MMP, we have
$A_{Y,f^{-1}_{*}\Delta+E+f^{*}\pi^{*}H}(F)\geq A_{Y^{\prime},f^{\prime-1}_{*}\Delta+\phi_{*}E+f^{\prime*}\pi^{*}H}(F)$
for every prime divisor $F$, and the equality holds if and only if $\phi$ is a local isomorphism at $\mathrm{center}_{Y}(F)$. Thus $\phi$ is a local isomorphism at every lc center of $(Y,f^{-1}_{*}\Delta+E+f^{*}\pi^{*}H)$. In particular, $\phi$ does not contract any component $E_{i}$ of $E$. It follows that $f^{\prime}\colon(Y^{\prime},E^{\prime}=\phi_{*}E)\to(X,\Delta)$ is the desired $\mathbb{Q}$-factorial semi-ample model satisfying (1), (2), and (3).
It remains to show (4). Take a model as above such that $Y$ is $\mathbb{Q}$-factorial and $-(K_{Y}+f^{-1}_{*}\Delta+E)$ is $f$-semi-ample. As $X$ is affine, let
$G\in|{-(K_{Y}+f^{-1}_{*}\Delta+E)}|_{\mathbb{Q}}$
be a general $\mathbb{Q}$-divisor, such that $(Y,f^{-1}_{*}\Delta+G+E+f^{*}\pi^{*}H)$ is lc. Let $\Gamma=f_{*}G\sim_{\mathbb{Q}}-(K_{X}+\Delta)$, then
$f^{*}(K_{X}+\Delta+\Gamma+\pi^{*}H)=K_{Y}+f^{-1}_{*}\Delta+G+E+f^{*}\pi^{*}H,$
so $\pi\colon(X,\Delta+\Gamma)\to S$ is locally stable, and each $E_{i}$ is an lc place of $(X,\Delta+\Gamma)$. Since the function
$v\mapsto A_{X,\Delta+\Gamma}(v)=A_{X,\Delta}(v)-v(\Gamma)$
is non-negative, homogeneous, and convex on $\tau$, while vanishing at $\mathrm{ord}_{E_{i}}$, we get $A_{X,\Delta+\Gamma}(v)=0$ for all $v\in\tau$, that is, $\tau\subseteq\mathrm{LCP}(X,\Delta+\Gamma)$. ∎
Remark.
Assume that $v_{0}$ is a special valuation, and $\sigma$ is a special quasi-monomial cone. Then $\tau$ is also special, so $f_{\eta}\colon(Y_{\eta},E_{\eta})\to(X_{\eta},\Delta_{\eta})$ is a model of qdlt Fano type by Lemma 3.3, with $\mathrm{QM}(Y_{\eta},E_{\eta})=\tau$. In particular, $v_{0}\in\mathrm{QM}(Y_{\eta},E_{\eta})$.
In this case, none of $E_{i}$ is contracted in the ample model for $-(K_{Y}+f^{-1}_{*}\Delta+E)$ over $X$, since on the generic fiber we get the Kollár model for $\tau$. Thus we may assume that $-(K_{Y}+f^{-1}_{*}\Delta+E)$ is ample as well. Since $(Y_{\eta},(f_{\eta})^{-1}_{*}\Delta_{\eta}+E_{\eta})$ is qdlt, and every lc center of $(Y,f^{-1}_{*}\Delta+E)$ lies over $\eta$ (see [Kol_fam, Cor. 4.56]), we conclude that $(Y,f^{-1}_{*}\Delta+E)$ is qdlt. However, $f\colon(Y,E)\to(X,\Delta)$ need not be a locally stable family of Kollár models over $S$.
Definition 4.3.
Let $\pi\colon(X,\Delta)\to S$ be a family of pairs, and $f\colon(Y,E)\to(X,\Delta)$ be a family of models over $S$ where $E=\sum_{i=1}^{k}E_{i}$ is the sum of all $f$-exceptional divisors. Let $Z$ be an irreducible component of $E_{I}=\bigcap_{i\in I}E_{i}$. Suppose $\zeta$ is the generic point of $Z$, and $z$ is a generic point of $Z_{s}=Z\cap Y_{s}$. Assume that $\pi\circ f\colon(Y,f^{-1}_{*}\Delta+E)\to S$ is locally stable in an open neighborhood of $z$, and each $E_{i}$ is $\mathbb{Q}$-Cartier at $z$. Then $(Y,f^{-1}_{*}\Delta+E+f^{*}\pi^{*}H)$ is lc near $z$, and $Z$ is an lc center thereof. Thus its generic point $\zeta$ lies over $\eta$. Since each $E_{i,\eta}$ is $\mathbb{Q}$-Cartier, $(Y_{\eta},E_{\eta})$ is simple-toroidal at $\zeta$ by Lemma 2.2. Now
$z\in\bigcap_{i\in I}E_{i}\cap\bigcap_{j=1}^{d}f^{*}\pi^{*}H_{j}$
where each $E_{i}$ is $\mathbb{Q}$-Cartier, and each $f^{*}\pi^{*}H_{j}$ is Cartier. Since $z$ has codimension at most $|I|+d$ in $Y$, $(Y,E+f^{*}\pi^{*}H)$ is simple-toroidal at $z$ by Lemma 2.2, hence $(Y_{s},E_{s})$ is simple-toroidal at $z$ as well. By Lemma 2.8, there are canonical maps
[…]
which sends $E_{i,\eta}$ to the unique irreducible component $E_{i,s}^{\prime}$ of $E_{i,s}$ containing $z$.
For $v^{\alpha}_{\eta}\in\mathrm{QM}_{\zeta}(Y_{\eta},E_{\eta})$, its image $v^{\alpha}_{s}\in\mathrm{QM}_{z}(Y_{s},E_{s})$ is called a quasi-monomial degeneration of $v^{\alpha}_{\eta}$ via the family of models $f\colon(Y,E)\to(X,\Delta)$, and $v^{\alpha}_{\eta}$ is said to degenerates to $v^{\alpha}_{s}$ via $(Y,E)$.
Lemma 4.4.
Let $\pi\colon(X,\Delta)\to S$ be a family of pairs. Suppose $v_{\eta}\in\mathrm{Val}_{X_{\eta}}^{\mathrm{qm}}$ degenerates to $v_{s}\in\mathrm{Val}_{X_{s}}^{\mathrm{qm}}$ via a family of models $f\colon(Y,E=\sum_{i=1}^{k}E_{i})\to(X,\Delta)$ over $S$. Then the following hold:
  1. (1)
    $v_{\eta}(\mathfrak{a}_{\eta})\leq v_{s}(\mathfrak{a}_{s})$ for every coherent ideal $\mathfrak{a}\subseteq\mathscr{O}_{X}$;
  2. (2)
    if $K_{X}+\Delta$ is $\mathbb{Q}$-Cartier, then $A_{X_{\eta},\Delta_{\eta}}(v_{\eta})=A_{X_{s},\Delta_{s}}(v_{s})$;
  3. (3)
    if $\Gamma$ is a relative $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisor on $X/S$ such that $\pi\colon(X,\Delta+\Gamma)\to S$ is locally stable and $v_{\eta}\in\mathrm{LCP}(X_{\eta},\Delta_{\eta}+\Gamma_{\eta})$, then $v_{s}\in\mathrm{LCP}(X_{s},\Delta_{s}+\Gamma_{s})$ and $v_{\eta}(\Gamma_{\eta})=v_{s}(\Gamma_{s})$.
Proof.
Suppose $Z$ is an irreducible component $\bigcap_{i\in I}E_{i}$ as in Definition 4.3, such that $v_{\eta}$ maps to $v_{s}$ in
$\mathrm{QM}_{\zeta}(Y_{\eta},E_{\eta})\simeq\mathbb{R}_{\geq 0}^{r}\simeq\mathrm{QM}_{z}(Y_{s},E_{s}),$
where $\zeta$ is the generic point of $Z_{\eta}$, and $z$ is a generic point of $Z_{s}$, and they are given by $\alpha=(\alpha_{i})\in\mathbb{R}_{\geq 0}^{I}$ in coordinates. Then (1) follows from Lemma 2.8.
(2). Write
$f^{*}(K_{X}+\Delta+\pi^{*}H)=K_{Y}+f^{-1}_{*}\Delta+\sum_{i=1}^{k}(1-A_{X,\Delta}(E_{i}))\cdot E_{i}+f^{*}\pi^{*}H.$
Restricting to $f_{\eta}\colon(Y_{\eta},\Delta_{\eta})\to(X_{\eta},\Delta_{\eta})$ and $f_{s}\colon(Y_{s},\Delta_{s})\to(X_{s},\Delta_{s})$, respectively, by adjunction we get
$A_{X_{\eta},\Delta_{\eta}}(E_{i,\eta})=A_{X,\Delta}(E_{i})=A_{X_{s},\Delta_{s}}(E_{i,s}^{\prime})$
for every component $E_{i,s}^{\prime}$ of $E_{i,s}$. Hence $A_{X_{\eta},\Delta_{\eta}}(v_{\eta})=\sum_{i\in I}\alpha_{i}A_{X,\Delta}(E_{i})=A_{X_{s},\Delta_{s}}(v_{s})$.
(3). By (1) and (2) we have
$A_{X_{s},\Delta_{s}+\Gamma_{s}}(v_{s})=A_{X_{s},\Delta_{s}}(v_{s})-v_{s}(\Gamma_{s})\leq A_{X_{\eta},\Delta_{\eta}}(v_{\eta})-v_{\eta}(\Gamma_{\eta})=A_{X_{\eta},\Delta_{\eta}+\Gamma_{\eta}}(v_{\eta})=0,$
while $A_{X_{s},\Delta_{s}+\Gamma_{s}}(v_{s})\geq 0$ since $(X_{s},\Delta_{s}+\Gamma_{s})$ is slc. Thus all the inequalities above must be equalities, that is, $A_{X_{s},\Delta_{s}+\Gamma_{s}}(v_{s})=0$ and $v_{s}(\Gamma_{s})=v_{\eta}(\Gamma_{\eta})$. ∎
Lemma 4.5.
Let $\pi\colon x\in(X,\Delta)\to S$ be a locally stable family of klt singularities, and $I_{\bullet}\subseteq\mathscr{O}_{X}$ be an ideal sequence on $X$ cosupported on $\{x\}$ such that
$\mathrm{lct}(X_{s},\Delta_{s};I_{s,\bullet})=\mathrm{lct}(X_{\eta},\Delta_{\eta};I_{\eta,\bullet})<\infty.$
Suppose $v_{s}\in\mathrm{Val}_{X_{s},x_{s}}^{\mathrm{qm}}$ and $v_{\eta}\in\mathrm{Val}_{X_{\eta},x_{\eta}}^{\mathrm{qm}}$ are special valuations with
$A_{X_{s},\Delta_{s}}(v_{s})=A_{X_{\eta},\Delta_{\eta}}(v_{\eta})=1,$
satisfying the following: for every relative effective $\mathbb{Q}$-Cartier divisor $D$ on $X/S$ with $v_{s}(D_{s})=v_{\eta}(D_{\eta})$, there exists $\epsilon>0$ such that $\pi\colon(X,\Delta+\epsilon D)\to S$ has klt fibers, and $v_{s}$ and $v_{\eta}$ are, up to scaling, unique quasi-monomial valuations computing lc thresholds $\mathrm{lct}(X_{s},\Delta_{s}+\epsilon D_{s};I_{s,\bullet})$ and $\mathrm{lct}(X_{\eta},\Delta_{\eta}+\epsilon D_{\eta};I_{\eta,\bullet})$, respectively. Then there exists a locally stable family of Kollár models at $x$
$\bar{f}\colon(\overline{Y},\overline{E})\to(X,\Delta)$
such that $v_{s}\in\mathrm{QM}(\overline{Y}_{s},\overline{E}_{s})$, $v_{\eta}\in\mathrm{QM}(\overline{Y}_{\eta},\overline{E}_{\eta})$, and $v_{\eta}$ degenerates to $v_{s}$.
Proof.
First note that for $D=0$ we have
$v_{s}(I_{s,\bullet})=A_{X_{s},\Delta_{s}}(v_{s})\cdot\mathrm{lct}(X_{s},\Delta_{s};I_{s,\bullet})=A_{X_{\eta},\Delta_{\eta}}(v_{\eta})\cdot\mathrm{lct}(X_{\eta},\Delta_{\eta};I_{\eta,\bullet})=v_{\eta}(I_{\eta,\bullet}).$
Assume $D$ is a relative effective $\mathbb{Q}$-Cartier divisor on $X/S$ with $v_{s}(D_{s})=v_{\eta}(D_{\eta})$, then
$\mathrm{lct}(X_{s},\Delta_{s}+\epsilon D_{s};I_{s,\bullet})$ $=\frac{A_{X_{s},\Delta_{s}}(v_{s})-\epsilon v_{s}(D_{s})}{v_{s}(I_{s,\bullet})}$
$=\frac{A_{X_{\eta},\Delta_{\eta}}(v_{\eta})-\epsilon v_{\eta}(D_{\eta})}{v_{\eta}(I_{\eta,\bullet})}=\mathrm{lct}(X_{\eta},\Delta_{\eta}+\epsilon D_{\eta};I_{\eta,\bullet}).$
Thus we can apply Lemma 4.2 to the family $\pi\colon x\in(X,\Delta+\epsilon D)\to S$, the ideal sequence $I_{\bullet}$, and $v_{\eta}$, which gives a family of models $f^{D}\colon(Y^{D},E^{D})\to(X,\Delta)$, via which $v_{\eta}$ degenerates to some $v^{D}_{s}\in\mathrm{Val}_{X_{s},x_{s}}^{\mathrm{qm}}$ as in Definition 4.3 (see the remark after Lemma 4.2). Then
$\mathrm{lct}(X_{\eta},\Delta_{\eta}+\epsilon D_{\eta};I_{\eta,\bullet})=\frac{A_{X_{\eta},\Delta_{\eta}+\epsilon D_{\eta}}(v_{\eta})}{v_{\eta}(I_{\eta,\bullet})}\geq\frac{A_{X_{s},\Delta_{s}+\epsilon D_{s}}(v_{s}^{D})}{v_{s}^{D}(I_{s,\bullet})}$
by Lemma 4.4. This implies that $v_{s}^{D}$ computes $\mathrm{lct}(X_{s},\Delta_{s}+\epsilon D_{s};I_{s,\bullet})$. By the uniqueness assumption, we conclude that $v_{s}^{D}=v_{s}$ since $A_{X_{s},\Delta_{s}+\epsilon D_{s}}(v_{s}^{D})=A_{X_{\eta},\Delta_{\eta}+\epsilon D_{\eta}}(v_{\eta})=A_{X_{s},\Delta_{s}+\epsilon D_{s}}(v_{s})$. Moreover, there is a relative effective $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisor $\Gamma\geq\epsilon D$ on $X/S$ such that $\pi\colon(X,\Delta+\Gamma)\to S$ is locally stable, $K_{X}+\Delta+\Gamma\sim_{S,\mathbb{Q}}0$, and $v_{\eta}\in\mathrm{LCP}(X_{\eta},\Delta_{\eta}+\Gamma_{\eta})$. By Lemma 4.4, we get $v_{s}\in\mathrm{LCP}(X_{s},\Delta_{s}+\Gamma_{s})$.
Now we apply Lemma 4.2 for $D=0$, and choose a family of models $f\colon(Y,E=\sum_{i=1}^{r}E_{i})\to(X,\Delta)$ such that each $E_{i}$ is $\mathbb{Q}$-Cartier, and there exists an $f$-ample divisor supported on $E$. Moreover, we may assume that $Z=\bigcap_{i=1}^{r}E_{i}$ is irreducible, with the generic point $\zeta$ such that $v_{\eta}\in\mathrm{QM}_{\zeta}(Y_{\eta},E_{\eta})$. Let $z$ be a generic point of $Z_{s}$, then $(Y,E+f^{*}\pi^{*}H)$ is simple-toroidal at $z$, and $v_{s}\in\mathrm{QM}_{z}(Y_{s},E_{s})$ such that $v_{\eta}$ and $v_{s}$ corresponds to $\alpha=(\alpha_{i})\in\mathbb{R}_{\geq 0}^{r}$ under
$\mathrm{QM}_{\zeta}(Y_{\eta},E_{\eta})\simeq\mathbb{R}_{\geq 0}^{r}\simeq\mathrm{QM}_{z}(Y_{s},E_{s}).$
In fact, by the uniqueness assumption, $z$ is the unique generic point of $Z_{s}$. Let
$\mu\colon(W,F)\to(Y,f^{-1}_{*}\Delta+E+f^{*}\pi^{*}H)$
be a log resolution that is a local isomorphism over $z$ and the generic point of $Y_{s}$, such that there is a $\mu$-ample divisor supported on $\mathrm{Ex}(\mu)$. Hence there is a $g$-ample divisor $-A$ with $\mathrm{Supp}(A)=\mathrm{Ex}(g)$ and $A\geq 0$, where $g=f\circ\mu$. Since $X$ is affine, $-A$ is ample. Let $G\in|{-A}|_{\mathbb{Q}}$ be a general $\mathbb{Q}$-divisor whose support does not contain any stratum of $(W,F)$, and let $D=g_{*}G=g_{*}(G+A)$. Then $D$ is $\mathbb{Q}$-Cartier, and $\mathrm{Supp}(D)$ does not contain $X_{s}$. Thus $D$ is flat over $S$. Write
$f^{*}D=f^{-1}_{*}D+\sum_{i=1}^{r}\mathrm{ord}_{E_{i}}(D)E_{i}$
where the support of $f^{-1}_{*}D=\mu_{*}G$ does not contain $z$. Hence $v_{s}(D_{s})=\sum_{i=1}^{r}\alpha_{i}\mathop{\mathrm{ord}_{E_{i}}}(D)=v_{\eta}(D_{\eta})$. By the argument above, there exists a relative effective $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisor $\Gamma\geq\epsilon D$ on $X/S$, for some $\epsilon>0$, such that $\pi\colon(X,\Delta+\Gamma)\to S$ is locally stable, $K_{X}+\Delta+\Gamma\sim_{S,\mathbb{Q}}0$, and $v_{\eta}\in\mathrm{LCP}(X_{\eta},\Delta_{\eta}+\Gamma_{\eta})$.
Moreover, consider the diagram
[…]
Suppose $\alpha=(\alpha_{i})\in\mathbb{R}_{\geq 0}^{r}$ corresponds to $v_{\eta}^{\alpha}\in\mathrm{QM}_{\zeta}(Y_{\eta},E_{\eta})\cap\mathrm{LCP}(X_{\eta},\Delta_{\eta}+\Gamma_{\eta})$, then
$v_{s}^{\alpha}\in\mathrm{QM}_{z}(Y_{s},E_{s})\cap\mathrm{LCP}(X_{s},\Delta_{s}+\Gamma_{s})$
and $v_{s}^{\alpha}(\Gamma_{s})=v_{\eta}^{\alpha}(\Gamma_{\eta})$ by Lemma 4.4. For every $\beta=(\beta_{j})\in\mathbb{R}_{\geq 0}^{d}$ and $v^{\alpha,\beta}\in\mathrm{QM}_{z}(Y,E+f^{*}\pi^{*}H)$,
$A_{X,\Delta+\pi^{*}H}(v^{\alpha,\beta})=\sum_{i=1}^{r}\alpha_{i}A_{X,\Delta+\pi^{*}H}(E_{i})+\sum_{j=1}^{d}\beta_{j}A_{X,\Delta+\pi^{*}H}(f^{*}\pi^{*}H_{j})=\sum_{i=1}^{r}\alpha_{i}A_{X,\Delta+\pi^{*}H}(E_{i})$
since $A_{X,\Delta+\pi^{*}H}(f^{*}\pi^{*}H_{j})=0$. Thus $A_{X_{\eta},\Delta_{\eta}}(v^{\alpha}_{\eta})=A_{X,\Delta+\pi^{*}H}(v^{\alpha,\beta})=A_{X_{s},\Delta_{s}}(v_{s}^{\alpha})$. Also,
$v_{\eta}^{\alpha}(\Gamma_{\eta})\leq v^{\alpha,\beta}(\Gamma)\leq v_{s}^{\alpha}(\Gamma_{s})$
by Lemma 2.8, so all the equality holds. Then we get
$A_{X_{\eta},\Delta_{\eta}+\Gamma_{\eta}}(v^{\alpha}_{\eta})=A_{X,\Delta+\Gamma+\pi^{*}H}(v^{\alpha,\beta})=A_{X_{s},\Delta_{s}+\Gamma_{s}}(v_{s}^{\alpha})=0.$
Let $\sigma_{\eta}\subseteq\mathrm{QM}_{\zeta}(Y_{\eta},E_{\eta})\cap\mathrm{LCP}(X_{\eta},\Delta_{\eta}+\Gamma_{\eta})$ be a simplicial cone containing $v_{\eta}$, and let $\sigma\subseteq\mathbb{R}_{\geq 0}^{r}$ be the corresponding cone, then the cone $\Sigma\subseteq\mathrm{QM}_{z}(Y,E+f^{*}\pi^{*}H)$ corresponding to $\sigma\times\mathbb{R}_{\geq 0}^{d}$ satisfies
$\Sigma\subseteq\mathrm{QM}_{z}(Y,E+f^{*}\pi^{*}H)\cap\mathrm{LCP}(X,\Delta+\Gamma+\pi^{*}H).$
Now $\Gamma+\pi^{*}H$ is a special $\mathbb{Q}$-complement with respect to $g\colon(W,F)\to(X,\Delta)$, since $g^{-1}_{*}(\Gamma+\pi^{*}H)\geq\epsilon G$. Thus, by Lemma 3.7, we get a qdlt anti-canonical model
$\bar{f}\colon(\overline{Y},\overline{E}+\overline{H})\to(X,\Delta)$
with $\mathrm{QM}(\overline{Y},\overline{E}+\overline{H})=\Sigma$. Thus all components of $\overline{E}$ are centered at $x$, and $\overline{H}=\bar{f}^{-1}_{*}\pi^{*}H=\bar{f}^{*}\pi^{*}H$. So $\bar{f}\colon(\overline{Y},\overline{E})\to(X,\Delta)$ is a locally stable family of Kollár models (see the remark after Definition 3.4). It is clear that $v_{\eta}\in\mathrm{QM}(\overline{Y}_{\eta},\overline{E}_{\eta})$ degenerates to $v_{s}\in\mathrm{QM}(\overline{Y}_{s},\overline{E}_{s})$. ∎
4.2. Multiple equivariant degenerations
Definition 4.6.
Let $x\in(X,\Delta)$ be a klt singularity of finite type over a field $k$, with $x\in X(k)$. Suppose
$f\colon(Y,E=E_{1}+\cdots+E_{r})\to(X,\Delta)$
is a model of qdlt Fano type at $x$. Let $\mathcal{R}$ be the extended Rees algebra
$\mathcal{R}\coloneqq\bigoplus_{(a_{1},\ldots,a_{r})\in\mathbb{Z}^{r}}f_{*}\mathscr{O}_{Y}(a_{1}E_{1}+\cdots+a_{r}E_{r})t_{1}^{a_{1}}\cdots t_{r}^{a_{r}}\subseteq\mathscr{O}_{X}[t_{1}^{\pm 1},\ldots,t_{r}^{\pm 1}].$
Then $\mathcal{R}$ is a finitely generated $k$-algebra, and we have $\mathcal{X}=\mathop{\mathrm{Spec}}(\mathcal{R})\to\mathbb{A}^{r}_{k}$ with an isomorphism
$\mathcal{X}\times_{\mathbb{A}^{r}_{k}}(\mathbb{G}_{\mathrm{m},k})^{r}\simeq X\times_{k}(\mathbb{G}_{\mathrm{m},k})^{r}.$
Let $\Delta_{\mathcal{X}}$ be the closure of $\Delta\times_{k}(\mathbb{G}_{\mathrm{m},k})^{r}$. The morphism $(\mathcal{X},\Delta_{\mathcal{X}})\to\mathbb{A}^{r}_{k}$ is called the multiple degeneration of $(X,\Delta)$ induced by $f\colon(Y,E)\to(X,\Delta)$. Moreover, there is an action of $(\mathbb{G}_{\mathrm{m},k})^{r}$ on $(\mathcal{X},\Delta_{\mathcal{X}})$ such that the multiple degeneration is equivariant, and the point $x\in X(k)$ induces a section $x\in\mathcal{X}(\mathbb{A}^{r}_{k})$.
By [XZ_stable_deg, §4], $x\in(\mathcal{X},\Delta_{\mathcal{X}})\to\mathbb{A}_{k}^{r}$ is a locally stable family of klt singularities, and
$\mathcal{R}/(t_{1},\ldots,t_{r})\simeq\mathrm{gr}_{v}(\mathscr{O}_{X})$
for every $v\in\mathrm{QM}^{\circ}(Y,E)$, which induces $(X_{v},\Delta_{v})\simeq(\mathcal{X},\Delta_{\mathcal{X}})\times_{\mathbb{A}^{r}_{k}}\{0\}$.
Lemma 4.7.
Let $S$ be a regular scheme. Suppose $\pi\colon x\in(X,\Delta)\to S$ is a locally stable family of klt singularities, and $f\colon(Y,E=E_{1}+\cdots+E_{r})\to(X,\Delta)$ is a locally stable family of Kollár models at $x$. For each $1\leq i\leq r$, let $(\mathcal{X}^{(i)},\Delta^{(i)})\to\mathbb{A}^{i}_{S}$ be the multiple degeneration induced by $E_{1}+\cdots+E_{i}$, that is,
$\mathcal{R}^{(i)}\coloneqq\bigoplus_{(a_{1},\ldots,a_{i})\in\mathbb{Z}^{i}}\pi_{*}f_{*}\mathscr{O}_{Y}(a_{1}E_{1}+\cdots+a_{i}E_{i})t_{1}^{a_{1}}\cdots t_{i}^{a_{i}}\subseteq\pi_{*}\mathscr{O}_{X}[t_{1}^{\pm 1},\ldots,t_{i}^{\pm 1}],$
and $\mathcal{X}^{(i)}=\mathop{\mathrm{Spec}_{S}}\mathcal{R}^{(i)}\to\mathbb{A}^{i}_{S}$. Then the following hold:
  1. (1)
    $(\mathcal{X}^{(i)},\Delta^{(i)})\to\mathbb{A}^{i}_{S}$ is a locally stable family of klt pairs.
  2. (2)
    There exists a locally stable family of Kollár models $f^{(i)}\colon(\mathcal{Y}^{(i)},\mathcal{E}^{(i)})\to(\mathcal{X}^{(i)},\Delta^{(i)})$ over $\mathbb{A}^{i}_{S}$ with
    $(\mathcal{Y}^{(i)},\mathcal{E}^{(i)})\times_{\mathbb{A}^{i}_{S}}(\mathbb{G}_{\mathrm{m},S})^{i}\simeq(Y,E)\times_{S}(\mathbb{G}_{\mathrm{m},S})^{i}$
    as families of models for $(X,\Delta)\times_{S}(\mathbb{G}_{\mathrm{m},S})^{i}$.
  3. (3)
    $(\mathcal{X}^{(i+1)},\Delta^{(i+1)})\to\mathbb{A}^{i+1}_{S}$ is the degeneration of $(\mathcal{X}^{(i)},\Delta^{(i)})\to\mathbb{A}^{i}_{S}$ induced by the component $\mathcal{E}^{(i)}_{i+1}$ corresponding to $E_{i+1}$.
  4. (4)
    For every $s\in S$, the base change $(\mathcal{X}^{(i)}_{s},\Delta^{(i)}_{s})\to\mathbb{A}^{i}_{s}$ is the multiple degeneration of $x_{s}\in(X_{s},\Delta_{s})$ induced by $f_{s}\colon(Y_{s},E_{1,s}+\cdots+E_{i,s})$.
Proof.
The case when $S$ is a point is proved in [XZ_stable_deg, §4.2]. In general, we may assume that $S$ is affine. If we have (1) and (2), then (3) holds by [Xu_towards, Prop. 3.6]. We prove (1) and (2) by induction on $i$. By (3), it suffices to prove the case $i=1$. By shrinking $S$, we may assume that there is a relative effective $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisor $\Gamma$ for $X/S$ such that $\pi\colon(X,\Delta+\Gamma)\to S$ is locally stable, $K_{X}+\Delta+\Gamma\sim_{S,\mathbb{Q}}0$, and
$\mathrm{LCP}(X_{s},\Delta_{s}+\Gamma_{s})=\mathrm{QM}(Y_{s},E_{s})$
for every $s\in S$. Now
$\mathcal{X}^{(1)}=\mathop{\mathrm{Spec}_{S}}\mathopen{}\mathclose{{}\left(\bigoplus_{m\in\mathbb{Z}}\pi_{*}f_{*}\mathscr{O}_{Y}(mE_{1})t^{m}}\right)\to\mathbb{A}^{1}_{S}=\mathop{\mathrm{Spec}_{S}}(\mathscr{O}_{S}[t]).$
By [XZ_stable_deg, Lem. 3.13(4)], $E_{1}$ gives a locally stable family of Kollár components for $(X,\Delta+(1-\epsilon)\Gamma)$ at $x$ for all $0<\epsilon\leq 1$, and the fiber $\mathcal{X}^{(1)}_{s}\to\mathbb{A}^{1}_{s}$ is the degeneration of $X_{s}$ induced by $E_{1,s}$ for every $s\in S$. Hence by [Xu_towards, Lem. 3.3] and the case over a point, $(\mathcal{X}^{(1)},\Delta^{(1)}+(1-\epsilon)\Gamma^{(1)})\to\mathbb{A}^{1}_{S}$ is a locally stable family with klt fibers. It follows that $(\mathcal{X}^{(1)},\Delta^{(1)}+\Gamma^{(1)})\to\mathbb{A}^{1}_{S}$ is a locally stable family, and (4) holds.
It remains to prove (2) when $i=1$. Since $(\mathcal{X}^{(1)},\Delta^{(1)}+\Gamma^{(1)})\times_{\mathbb{A}^{1}_{S}}\mathbb{G}_{\mathrm{m},S}\simeq(X,\Delta+\Gamma)\times_{S}\mathbb{G}_{\mathrm{m},S}$, we can extract the divisors $E_{i}\times_{S}\mathbb{G}_{\mathrm{m},S}$ to get a model
$f^{(1)}\colon(\mathcal{Y}^{(1)},\mathcal{E}^{(1)})\to(\mathcal{X}^{(1)},\Delta^{(1)}).$
by Lemma 2.4. Passing to the ample model over $\mathcal{X}^{(1)}$, we may assume $-(K_{\mathcal{Y}^{(1)}}+(f^{(1)})^{-1}_{*}\Delta^{(1)}+\mathcal{E}^{(1)})$ is ample, and there is an isomorphism
$(\mathcal{Y}^{(1)},\mathcal{E}^{(1)})\times_{\mathbb{A}^{1}_{S}}\mathbb{G}_{\mathrm{m},S}\simeq(Y,E)\times_{S}\mathbb{G}_{\mathrm{m},S}.$
Hence $\mathcal{E}^{(1)}$ is the closure of $E\times_{S}\mathbb{G}_{\mathrm{m},S}$, and
$K_{\mathcal{Y}^{(1)}}+(f^{(1)})^{-1}_{*}(\Delta^{(1)}+\Gamma^{(1)})+\mathcal{E}^{(1)}=(f^{(1)})^{*}(K_{\mathcal{X}^{(1)}}+\Delta^{(1)}+\Gamma^{(1)}).$
Suppose $t_{0},\ldots,t_{d}$ is a regular system of parameters at a closed point of $\mathbb{A}^{1}_{S}$, and
$H=\mathop{\mathrm{div}}(t_{0})+\cdots+\mathop{\mathrm{div}}(t_{d})\subseteq\mathcal{X}^{(1)}.$
Then $(\mathcal{X}^{(1)},\Delta^{(1)}+\Gamma^{(1)}+H)$ is lc. Hence $(\mathcal{Y}^{(1)},\mathcal{E}^{(1)}+(f^{(1)})^{-1}_{*}(\Delta^{(1)}+\Gamma^{(1)})+(f^{(1)})^{*}H)$ is lc, that is,
$(\mathcal{Y}^{(1)},\mathcal{E}^{(1)}+(f^{(1)})^{-1}_{*}(\Delta^{(1)}+\Gamma^{(1)}))\to\mathbb{A}^{1}_{S}$
is locally stable. Hence, for every $s\in S$, the fiber $f_{s}^{(1)}\colon(\mathcal{Y}^{(1)}_{s},\mathcal{E}^{(1)}_{s})\to(\mathcal{X}^{(1)}_{s},\Delta^{(1)}_{s})$ is a family of models which extends $(Y_{s},E_{s})\times_{\kappa(s)}\mathbb{G}_{\mathrm{m},\kappa(s)}\to(X_{s},\Delta_{s})\times_{\kappa(s)}\mathbb{G}_{\mathrm{m},\kappa(s)}$ such that $-(K_{\mathcal{Y}^{(1)}_{s}}+(f_{s}^{(1)})^{-1}_{*}\Delta_{s}^{(1)}+\mathcal{E}_{s}^{(1)})$ is ample. By the case over a point, $(Y_{s},E_{s})\times_{\kappa(s)}\mathbb{G}_{\mathrm{m},\kappa(s)}\to(X_{s},\Delta_{s})\times_{\kappa(s)}\mathbb{G}_{\mathrm{m},\kappa(s)}$ can also be extended to a locally stable family of Kollár models, which must coincide with $f_{s}^{(1)}\colon(\mathcal{Y}^{(1)}_{s},\mathcal{E}^{(1)}_{s})\to(\mathcal{X}^{(1)}_{s},\Delta^{(1)}_{s})$. So every fiber of $(\mathcal{Y}^{(1)},\mathcal{E}^{(1)}+(f^{(1)})^{-1}_{*}\Delta^{(1)})\to\mathbb{A}^{1}_{S}$ is qdlt, that is,
$f^{(1)}\colon(\mathcal{Y}^{(1)},\mathcal{E}^{(1)})\to(\mathcal{X}^{(1)},\Delta^{(1)})$
is a locally stable family of Kollár models. ∎
Lemma 4.8.
Let $S$ be a regular scheme. Suppose $\pi\colon x\in(X,\Delta)\to S$ is a locally stable family of klt singularities, and $f\colon(Y,E=\sum_{i=1}^{r}E_{i})\to(X,\Delta)$ is a locally stable family of Kollár models at $x$. Suppose $\alpha=(\alpha_{i})\in\mathbb{R}_{\geq 0}^{r}$, and $v_{s}^{\alpha}\in\mathrm{QM}(Y_{s},E_{s})$ is the corresponding quasi-monomial valuation for every $s\in S$. For each $\lambda\geq 0$, let
$\mathfrak{a}_{\lambda}\coloneqq\sum_{\langle\alpha,m\rangle\geq\lambda}f_{*}\mathscr{O}_{Y}\mathopen{}\mathclose{{}\left(-\sum_{i=1}^{r}m_{i}E_{i}}\right)\subseteq\mathscr{O}_{X},$
where the sum ranges over all $m=(m_{i})\in\mathbb{N}^{r}$ such that $\sum_{i=1}^{r}\alpha_{i}m_{i}\geq\lambda$. Then the following hold:
  1. (1)
    $\mathfrak{a}_{\bullet}=\{\mathfrak{a}_{\lambda}\}_{\lambda}$ is an ideal sequence on $X$ centered at $x$, and $\mathfrak{a}_{\lambda}/\mathfrak{a}_{>\lambda}$ is flat over $S$ for all $\lambda\geq 0$.
  2. (2)
    Let
    $X_{\alpha}=\mathop{\mathrm{Spec}_{S}}\bigoplus_{\lambda\geq 0}\mathfrak{a}_{\lambda}/\mathfrak{a}_{>\lambda}$
    with the canonical section $x_{\alpha}\in X_{\alpha}(S)$ and the induced divisor $\Delta_{\alpha}$, then $x_{\alpha}\in(X_{\alpha},\Delta_{\alpha})\to S$ is a locally stable family of klt singularities.
  3. (3)
    For every $s\in S$, the restriction $\mathfrak{a}_{s,\bullet}=\{\mathfrak{a}_{\lambda}\mathscr{O}_{X_{s}}\}_{\lambda}$ is the ideal sequence associated with $v_{s}^{\alpha}$.
  4. (4)
    For every $s\in S$, the fiber $x_{\alpha,s}\in(X_{\alpha,s},\Delta_{\alpha,s})$ is the degeneration of $x_{s}\in(X_{s},\Delta_{s})$ induced by $v^{\alpha}_{s}$.
Proof.
(1). It is clear that $\mathfrak{a}_{\bullet}=\{\mathfrak{a}_{\lambda}\}_{\lambda}$ is an ideal sequence on $X$. Write
$I_{m}=f_{*}\mathscr{O}_{Y}\mathopen{}\mathclose{{}\left(-\sum_{i=1}^{r}m_{i}E_{i}}\right)\subseteq\mathscr{O}_{X}$
for $m\in\mathbb{Z}^{r}$, and $I_{>m}\coloneqq\sum_{m^{\prime}>m}I_{m^{\prime}}$, where $m^{\prime}=(m^{\prime}_{i})>(m_{i})=m$ if and only if $m^{\prime}_{i}\geq m_{i}$ for all $i$ and $m^{\prime}\neq m$. By Lemma 4.7,
$\mathcal{R}\coloneqq\bigoplus_{m\in\mathbb{Z}^{r}}I_{m}t_{1}^{-m_{1}}\cdots t_{r}^{-m_{r}}.$
is flat over $\mathscr{O}_{S}[t_{1},\ldots,t_{r}]$. Hence
$\mathcal{R}/(t_{1},\ldots,t_{r})\simeq\bigoplus_{m\in\mathbb{Z}^{r}}I_{m}/I_{>m}$
is flat over $\mathscr{O}_{S}$. Note that $I_{m}\cap I_{n}=I_{\max(m,n)}$, where $\max(m,n)=(\max(m_{i},n_{i}))_{i}$. Hence
$\mathfrak{a}_{\lambda}/\mathfrak{a}_{>\lambda}\simeq\bigoplus_{\langle\alpha,m\rangle=\lambda}I_{m}/I_{>m}$
is also flat over $\mathscr{O}_{S}$.
(2). By the computation above, $(X_{\alpha},\Delta_{\alpha})\to S$ is isomorphic to the central fiber $(\mathcal{X},\Delta_{\mathcal{X}})\times_{\mathbb{A}^{r}_{S}}0_{S}\to S$ of the multiple degeneration in Lemma 4.7. Hence it is a locally stable family of klt singularities.
(3) and (4). For every $s\in S$, we construct $I_{s,m}\subseteq\mathscr{O}_{X_{s}}$ and $\mathcal{R}_{s}\coloneqq\bigoplus_{m\in\mathbb{Z}^{r}}I_{s,m}t_{1}^{-m_{1}}\cdots t_{r}^{-m_{r}}$ as above from $f_{s}\colon(Y_{s},E_{s})\to(X_{s},\Delta_{s})$. Then $\mathcal{R}_{s}=\mathcal{R}\otimes_{\mathscr{O}_{S}}\kappa(s)$ as subrings of $\mathscr{O}_{X_{s}}[t_{1}^{\pm 1},\ldots,t_{r}^{\pm 1}]$, by Lemma 4.7. So $I_{s,m}=I_{m}\mathscr{O}_{X_{s}}=I_{m}\otimes_{\mathscr{O}_{S}}\kappa(s)$ for all $m\in\mathbb{Z}^{r}$. By [XZ_stable_deg, Cor. 4.10], we have
$\mathfrak{a}_{\lambda}(v_{s}^{\alpha})=\sum_{\langle\alpha,m\rangle\geq\lambda}I_{s,m}$
Thus $\mathfrak{a}_{\lambda}(v_{s}^{\alpha})=\mathfrak{a}_{\lambda}\mathscr{O}_{X_{s}}$, and $\mathfrak{a}_{\lambda}(v_{s}^{\alpha})/\mathfrak{a}_{>\lambda}(v_{s}^{\alpha})\simeq(\mathfrak{a}_{\lambda}/\mathfrak{a}_{>\lambda})\otimes_{\mathscr{O}_{S}}\kappa(s)$. So we get (3) and (4). ∎
5. Families with constant local volume
5.1. Degeneration of the minimizer over a DVR
In this subsection, let $S=\mathop{\mathrm{Spec}}(A)$, where $A$ is a DVR, with the generic point $\eta\in S$ and the closed point $s\in S$.
Theorem 5.1.
Let $\pi\colon x\in(X,\Delta)\to S$ be a locally stable family of klt singularities with
$\widehat{\mathrm{vol}}(x_{\eta};X_{\eta},\Delta_{\eta})=\widehat{\mathrm{vol}}(x_{s};X_{s},\Delta_{s}).$
Suppose $v^{\mathrm{m}}_{\eta}\in\mathrm{Val}_{X_{\eta},x_{\eta}}$ and $v^{\mathrm{m}}_{s}\in\mathrm{Val}_{X_{s},x_{s}}$ are minimizers of the normalized volume for $x_{\eta}\in(X_{\eta},\Delta_{\eta})$ and $x_{s}\in(X_{s},\Delta_{s})$, respectively, scaled such that $A_{X_{\eta},\Delta_{\eta}}(v^{\mathrm{m}}_{\eta})=A_{X_{s},\Delta_{s}}(v^{\mathrm{m}}_{s})$. Then there exists a locally stable family of Kollár models $f\colon(Y,E)\to(X,\Delta)$ at $x$ over $S$ such that
$v^{\mathrm{m}}_{\eta}\in\mathrm{QM}(Y_{\eta},E_{\eta})\quad\text{and}\quad v^{\mathrm{m}}_{s}\in\mathrm{QM}(Y_{s},E_{s}),$
and they are identified under the canonical isomorphism $\mathrm{QM}(Y_{\eta},E_{\eta})\simeq\mathrm{QM}(Y_{s},E_{s})$.
Proof.
Let $\mathfrak{a}_{\bullet}(v^{\mathrm{m}}_{\eta})\subseteq\mathscr{O}_{X_{\eta}}$ and $\mathfrak{a}_{\bullet}(v^{\mathrm{m}}_{s})\subseteq\mathscr{O}_{X_{s}}$ be the ideal sequences associated with $v^{\mathrm{m}}_{\eta}$ and $v^{\mathrm{m}}_{s}$, respectively. We may assume that $A_{X_{\eta},\Delta_{\eta}}(v^{\mathrm{m}}_{\eta})=A_{X_{s},\Delta_{s}}(v^{\mathrm{m}}_{s})=1$, so that
$\mathrm{lct}(X_{\eta},\Delta_{\eta};\mathfrak{a}_{\bullet}(v^{\mathrm{m}}_{\eta}))=\mathrm{lct}(X_{s},\Delta_{s};\mathfrak{a}_{\bullet}(v^{\mathrm{m}}_{s}))=1,$
since they are computed by $v^{\mathrm{m}}_{\eta}$ and $v^{\mathrm{m}}_{s}$, respectively, by Lemma 2.17. Thus
$\mathrm{e}_{X_{\eta}}(\mathfrak{a}_{\bullet}(v^{\mathrm{m}}_{\eta}))=\widehat{\mathrm{vol}}(x_{\eta};X_{\eta},\Delta_{\eta})=\widehat{\mathrm{vol}}(x_{s};X_{s},\Delta_{s})=\mathrm{e}_{X_{s}}(\mathfrak{a}_{\bullet}(v^{\mathrm{m}}_{s})).$
Let $I_{\bullet}=j_{*}\mathfrak{a}_{\bullet}(v^{\mathrm{m}}_{\eta})\cap\mathscr{O}_{X}$, where $j\colon X_{\eta}\to X$ is the inclusion, and $\mathscr{O}_{X}\to j_{*}\mathscr{O}_{X_{\eta}}$ is injective since $X_{s}$ is a Cartier divisor. It is clear that $\mathscr{O}_{X}/I_{\lambda}$ is supported on the section $\{x\}$ and torsion-free over $S$ for all $\lambda>0$. Since $S$ is the spectrum of a DVR, we conclude that $\mathscr{O}_{X}/I_{\lambda}$ is flat over $S$. Hence
$\mathrm{e}_{X_{\eta}}(I_{\eta,\bullet})=\mathrm{e}_{X_{s}}(I_{s,\bullet}),$
where $I_{\eta,\bullet}=I_{\bullet}\mathscr{O}_{X_{\eta}}=\mathfrak{a}_{\bullet}(v^{\mathrm{m}}_{\eta})$, and $I_{s,\bullet}=I_{\bullet}\mathscr{O}_{X_{s}}$. Since $\mathrm{lct}$ is lower semi-continuous in a family, we have
$\widehat{\mathrm{vol}}(x_{s};X_{s},\Delta_{s})$ $\leq\mathrm{lct}(X_{s},\Delta_{s};I_{s,\bullet})^{n}\cdot\mathrm{e}_{X_{s}}(I_{s,\bullet})$
$\leq\mathrm{lct}(X_{\eta},\Delta_{\eta};I_{\eta,\bullet})^{n}\cdot\mathrm{e}_{X_{\eta}}(I_{\eta,\bullet})=\widehat{\mathrm{vol}}(x_{\eta};X_{\eta},\Delta_{\eta}),$
where $n=\dim_{x_{\eta}}(X_{\eta})=\dim_{x_{s}}(X_{s})$. But $\widehat{\mathrm{vol}}(x_{\eta};X_{\eta},\Delta_{\eta})=\widehat{\mathrm{vol}}(x_{s};X_{s},\Delta_{s})$, so all the inequalities above must be equalities. Thus by the uniqueness up to saturation in Lemma 2.17, we have
$I_{s,\bullet}\subseteq\widetilde{I}_{s,\bullet}=\mathfrak{a}_{c\bullet}(v^{\mathrm{m}}_{s})$
for some $c>0$. In fact, $c=1$ since $\mathrm{e}_{X_{s}}(\widetilde{I}_{s,\bullet})=\mathrm{e}_{X_{s}}(I_{s,\bullet})=\mathrm{e}_{X_{s}}(\mathfrak{a}_{\bullet}(v^{\mathrm{m}}_{s}))$.
Next, we will verify conditions of Lemma 4.5. Assume $D$ is a relative effective $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisor on $X/S$, such that
$v^{\mathrm{m}}_{\eta}(D_{\eta})=v^{\mathrm{m}}_{s}(D_{s}).$
By Lemma 3.8, there exists $\delta>0$ such that if $\epsilon<\delta/v^{\mathrm{m}}_{\eta}(D_{\eta})=\delta/v^{\mathrm{m}}_{s}(D_{s})$, then $v^{\mathrm{m}}_{\eta}$ and $v^{\mathrm{m}}_{s}$ are special valuations for $x_{\eta}\in(X_{\eta},\Delta_{\eta}+\epsilon D_{\eta})$ and $x_{s}\in(X_{s},\Delta_{s}+\epsilon D_{s})$, respectively. In fact, we can take $\delta=1/n$; see the remark after Lemma 3.8.
By Lemma 3.9, $v^{\mathrm{m}}_{\eta}$ and $v^{\mathrm{m}}_{s}$ are the unique, up to scaling, real valuations that compute lc thresholds $\mathrm{lct}(X_{\eta},\Delta_{\eta}+\epsilon D_{\eta};\mathfrak{a}_{\bullet}(v^{\mathrm{m}}_{\eta}))$ and $\mathrm{lct}(X_{s},\Delta_{s}+\epsilon D_{s};\mathfrak{a}_{\bullet}(v^{\mathrm{m}}_{s}))$, respectively. Then by Lemma 2.15, $v^{\mathrm{m}}_{s}$ is also the unique, up to scaling, real valuation that compute $\mathrm{lct}(X_{s},\Delta_{s}+\epsilon D_{s};I_{s,\bullet})$. Now the theorem follows from Lemma 4.5. ∎
5.2. General base schemes
Corollary 5.2.
Let $S$ be a semi-normal scheme. Let $\pi\colon x\in(X,\Delta)\to S$ be a locally stable family of klt singularities such that
$s\mapsto\widehat{\mathrm{vol}}(x_{s};X_{s},\Delta_{s})$
is a locally constant function on $S$. Suppose $v^{\mathrm{m}}_{s}\in\mathrm{Val}_{X_{s},x_{s}}$ is a minimizer of the normalized volume for $x_{s}\in(X_{s},\Delta_{s})$, scaled such that $A_{X_{s},\Delta_{s}}(v^{\mathrm{m}}_{s})=1$ for all $s\in S$. Then there is an ideal sequence $\mathfrak{a}_{\bullet}\subseteq\mathscr{O}_{X}$ centered at $x$, such that the following hold:
  1. (1)
    $\mathfrak{a}_{\lambda}/\mathfrak{a}_{>\lambda}$ is flat over $S$ for all $\lambda\geq 0$.
  2. (2)
    Let
    $X_{0}=\mathop{\mathrm{Spec}_{S}}\bigoplus_{\lambda\geq 0}\mathfrak{a}_{\lambda}/\mathfrak{a}_{>\lambda}$
    with the canonical section $x_{0}\in X_{0}(S)$ and the induced divisor $\Delta_{0}$, then $x_{0}\in(X_{0},\Delta_{0})\to S$ is a locally stable family of K-semistable log Fano cone singularities.
  3. (3)
    For every $s\in S$, $\mathfrak{a}_{s,\bullet}$ is the ideal sequence associated with $v_{s}^{\mathrm{m}}$, and the fiber $x_{0,s}\in(X_{0,s},\Delta_{0,s})$ is the K-semistable degeneration induced by $v_{s}^{\mathrm{m}}$.
Proof.
First assume that $S=\mathop{\mathrm{Spec}}(A)$, where $A$ is a DVR. By Theorem 5.1 and Lemma 4.8, we have an ideal sequence $\mathfrak{a}_{\bullet}\subseteq\mathscr{O}_{X}$ such that $\mathfrak{a}_{\lambda}/\mathfrak{a}_{>\lambda}$ is flat over $S$ for all $\lambda\geq 0$, $x_{0}\in(X_{0},\Delta_{0})\to S$ is a locally stable family of klt singularities, and $\mathfrak{a}_{s,\bullet}$ is the ideal sequence associated with $v_{s}^{\mathrm{m}}$ for each $s\in S$. Since $v_{s}^{\mathrm{m}}$ is a minimizer of the normalized volume function, $x_{0,s}\in(X_{0,s},\Delta_{0,s})$ is a K-semistable log Fano cone singularity by [LX_stability, Thm. 1.1].
In general, we may assume that $S$ is connected. Let $\mathfrak{a}_{\bullet}(v^{\mathrm{m}}_{s})\subseteq\mathscr{O}_{X_{s}}$ be the ideal sequence associated with $v^{\mathrm{m}}_{s}$. Note that if $\kappa(s)\subseteq K$ is field extension, then $\mathfrak{a}_{\bullet}(v^{\mathrm{m}}_{s})\otimes_{\kappa(s)}K$ is the ideal sequence associated with the minimizer of the normalized volume function for $x_{K}\in(X_{K},\Delta_{K})$. Suppose $A$ is a DVR, and $g\colon T=\mathop{\mathrm{Spec}}(A)\to S$ is a morphism, then the base change $x_{T}\in(X_{T},\Delta_{T})\to T$ is a locally stable family of klt singularities, so that we have an ideal sequence $\mathfrak{a}_{T,\bullet}\subseteq\mathscr{O}_{X_{T}}$, whose fibers are the base changes of the corresponding $\mathfrak{a}_{\bullet}(v^{\mathrm{m}}_{s})$. Therefore, the function
$\ell_{\lambda}\colon s\mapsto\mathrm{length}_{\mathscr{O}_{X_{s}}}\mathopen{}\mathclose{{}\left(\mathscr{O}_{X_{s}}/\mathfrak{a}_{\lambda}(v^{\mathrm{m}}_{s})}\right)$
is constant under specialization, hence constant, for all $\lambda\geq 0$. Consider the Hilbert scheme
$h_{\lambda}\colon\mathrm{Hilb}_{X/S}^{\ell_{\lambda}}\to S$
parameterizing closed subschemes of $X$ that are finite flat over $S$ with length $\ell_{\lambda}$. Then $h_{\lambda}$ is separated and of finite type. For each $s\in S$, we have a $\kappa(s)$-point $\sigma_{\lambda}(s)=[\mathscr{O}_{X_{s}}/\mathfrak{a}_{\lambda}(v^{\mathrm{m}}_{s})]\in h_{\lambda}^{-1}(s)$. If $A$ is a DVR, and $g\colon T=\mathop{\mathrm{Spec}}(A)\to S$ is a morphism, then by the argument above, there is a lifting $f\colon T\to\mathrm{Hilb}_{X/S}^{\ell_{\lambda}}$ of $g$ such that $f(t)=\sigma_{\lambda}(g(t))$ for every $t\in T$. Thus, by Lemma A.1, the set map $s\mapsto\sigma_{\lambda}(s)$ underlies a morphism of schemes $\sigma_{\lambda}\colon S\to\mathrm{Hilb}_{X/S}^{\ell_{\lambda}}$, which is a section of $h_{\lambda}$. Hence we get an ideal sequence $\mathfrak{a}_{\bullet}$ on $X$ such that $\mathfrak{a}_{s,\lambda}=\mathfrak{a}_{\lambda}(v^{\mathrm{m}}_{s})$ for all $s\in S$. Each $\mathfrak{a}_{\lambda}/\mathfrak{a}_{>\lambda}$ is a finitely generated $\mathscr{O}_{S}$-module with constant rank, hence flat over $S$.
Then $X_{0}=\mathop{\mathrm{Spec}_{S}}\bigoplus_{\lambda\geq 0}\mathfrak{a}_{\lambda}/\mathfrak{a}_{>\lambda}\to S$ is flat and of finite type, with normal and geometrically integral fibers $X_{0,s}=\mathop{\mathrm{Spec}}\bigoplus_{\lambda\geq 0}\mathfrak{a}_{\lambda}(v^{\mathrm{m}}_{s})/\mathfrak{a}_{>\lambda}(v^{\mathrm{m}}_{s})$. By a similar argument as above, using the representability of relative Mumford divisors (see [Kol_fam, Thm. 4.76]), there is a relative Mumford $\mathbb{Q}$-divisor $\Delta_{0}$ on $X_{0}/S$ such that $\Delta_{0,s}$ is the induced degeneration of $\Delta_{s}$ for all $s\in S$. Then $\Delta_{0}$ coincides with the degeneration of $\Delta$ defined by initial ideals, as they coincide on generic fibers. Finally, $x_{0}\in(X_{0},\Delta_{0})\to S$ is a locally stable family by the valuative criterion of local stability (see [Kol_fam, Thm. 4.7]). ∎
Corollary 5.3.
Let $S$ be a regular local scheme, and $\pi\colon x\in(X,\Delta)\to S$ be a locally stable family of klt singularities such that the function
$s\mapsto\widehat{\mathrm{vol}}(x_{s};X_{s},\Delta_{s})$
is constant on $S$. Suppose $v^{\mathrm{m}}_{s}\in\mathrm{Val}_{X_{s},x_{s}}$ is a minimizer of the normalized volume for $x_{s}\in(X_{s},\Delta_{s})$, scaled such that $A_{X_{s},\Delta_{s}}(v^{\mathrm{m}}_{s})=1$ for all $s\in S$. Then there exists a locally stable family of Kollár models $f\colon(Y,E=\sum_{i=1}^{r}E_{i})\to(X,\Delta)$ at $x$ over $S$ such that $v^{\mathrm{m}}_{s}\in\mathrm{QM}(Y_{s},E_{s})$, and the coordinate
$s\mapsto(v^{\mathrm{m}}_{s}(E_{i,s}))_{i}\in\mathbb{R}_{\geq 0}^{r}$
is a constant function on $S$.
Proof.
By Corollary 5.2, there is an ideal sequence $\mathfrak{a}_{\bullet}$ on $X$ centered at $x$ such that $\mathfrak{a}_{s,\bullet}=\mathfrak{a}_{\bullet}(v^{\mathrm{m}}_{s})$ for all $s\in S$. By the same argument as in the proof of Theorem 5.1, we can apply Lemma 4.5 to get the desired locally stable family of Kollár models. ∎
Corollary 5.4.
[…]
Proof.
The function
$s\mapsto\widehat{\mathrm{vol}}(x_{s};X_{s},\Delta_{s})$
is constructible and lower semi-continuous on $S$, by [Xu_qm, Thm. 1.3] and [BL_lsc, Thm. 1.1]; if the family degenerates to a family of K-semistable log Fano cone singularities, then the function is locally constant. Hence it suffices to take each $S_{i}\subseteq S$ to be a level set of this function by Corollary 5.2. ∎
Appendix A A valuative criterion for sections
Lemma A.1.
Let $S$ be a locally Noetherian semi-normal scheme, and $\pi\colon Z\to S$ be a morphism that is separated and of finite type. Suppose $|\sigma|\colon|S|\to|Z|$ is a map between underlying sets that is a section of the map $|\pi|\colon|Z|\to|S|$, such that for every morphism $g\colon T=\mathop{\mathrm{Spec}}(A)\to S$, where $A$ is a DVR, there exists a morphism $f\colon T\to Z$ such that $\pi\circ f=g$ and $|f|=|\sigma|\circ|g|$. Then there is a section $\sigma\colon S\to Z$ of $\pi$ whose underlying map between sets is $|\sigma|$.
Moreover, if $S$ is locally essentially of finite type over a field $k$, then it suffices to consider $A$ that are also essentially of finite type over $k$.
Proof.
We may assume $S$ is affine. If $T$ is the spectrum of a DVR, we write $\zeta\in T$ for the generic point, and $t\in T$ for the closed point. For $s\in S$ with residue field $\kappa=\kappa(s)$, let $A$ be a DVR with residue field $\kappa$, say, $A=\kappa[\varpi]_{(\varpi)}$, and consider the map
[…]
Then we conclude that $|\sigma|(s)\in|Z_{s}|$ is a $\kappa$-point of $Z_{s}$. In particular, it is a closed point of $|Z_{s}|$.
Let $|S^{\prime}|=|\sigma|(|S|)\subseteq|Z|$ be the image of $|\sigma|$. Note that $|\sigma|\colon|S|\to|S^{\prime}|$ is a bijection, and the restriction of $|\pi|$ is the inverse. Suppose $x^{\prime}\in|S^{\prime}|$ specializes to $y^{\prime}\in|Z|$, and $y^{\prime}\neq x^{\prime}$. Let $x=\pi(x^{\prime})$ and $y=\pi(y^{\prime})$. Since $x^{\prime}=|\sigma|(x)$ is a closed point in $Z_{x}$, we have $y\neq x$. Then there exists a DVR $A$ with fraction field $K=\kappa(x)$ and a morphism $f^{\prime}\colon T=\mathop{\mathrm{Spec}}(A)\to Z$ such that $f^{\prime}(\zeta)=x^{\prime}$, and $f^{\prime}(t)=y^{\prime}$ (see [stacks-project, Tag 00PH]), and we may assume that $A$ is essentially of finite type over $k$ if $S$ is so. By assumption, there exists a morphism $f\colon T\to Z$ such that $\pi\circ f=\pi\circ f^{\prime}$, $f(\zeta)=|\sigma|(x)=x^{\prime}$, and $f(t)=|\sigma|(y)$. Since $T$ has function field $K=\kappa(x)$, we must have $f|_{\mathop{\mathrm{Spec}}(K)}=f^{\prime}|_{\mathop{\mathrm{Spec}}(K)}$. Since $\pi$ is separated, we conclude that $f=f^{\prime}$ by the valuative criterion. Thus $y^{\prime}=f(t)=|\sigma|(y)\in|S^{\prime}|$. So $|S^{\prime}|$ is stable under specialization.
Let $\eta_{1},\ldots,\eta_{n}\in S$ be the generic points, and $\eta_{i}^{\prime}=|\sigma|(\eta_{i})\in|S^{\prime}|$. Then $\bigcup_{i=1}^{n}\overline{\{\eta_{i}^{\prime}\}}\subseteq|S^{\prime}|$. Conversely, if $x^{\prime}\in|S^{\prime}|$, then $x=\pi(x^{\prime})\in S$ generalizes to some $\eta_{i}$. Let $A$ be a DVR and $g\colon T=\mathop{\mathrm{Spec}}(A)\to S$ be a morphism such that $g(\zeta)=\eta_{i}$ and $g(t)=x$. By assumption, there is a lifting $f\colon T\to Z$ of $g$ such that $f(\zeta)=|\sigma|(\eta_{i})=\eta_{i}^{\prime}$ and $f(t)=|\sigma|(x)=x^{\prime}$. Hence $x^{\prime}\in\overline{\{\eta_{i}^{\prime}\}}$. Thus we conclude that $\bigcup_{i=1}^{n}\overline{\{\eta_{i}^{\prime}\}}\subseteq|S^{\prime}|$. Hence $|S^{\prime}|\subseteq|Z|$ is closed. Let $S^{\prime}\subseteq Z$ be the corresponding reduced closed subscheme.
By the valuative criterion, $\pi|_{S^{\prime}}\colon S^{\prime}\to S$ is proper. Moreover, it is bijective and induces isomorphisms on residue fields. Since $S$ is semi-normal, we conclude that $\pi|_{S^{\prime}}\colon S^{\prime}\to S$ is an isomorphism. Therefore we get a section $\sigma\colon S\simeq S^{\prime}\hookrightarrow Z$. ∎
[…]